### v1

# Long-term Causal Inference via Modeling Sequential Latent Confounding

Abstract
Long-term causal inference is an important but challenging problem across various scientific domains. To solve the latent confounding problem in long-term observational studies, existing methods leverage short-term experimental data. Ghassami et al. (2022) propose an approach based on the Conditional Additive Equi-Confounding Bias (CAECB) assumption, which asserts that the confounding bias in the short-term outcome is equal to that in the long-term outcome, so that the long-term confounding bias and the causal effects can be identified. While effective in certain cases, this assumption is limited to scenarios with a one-dimensional short-term outcome. In this paper, we introduce a novel assumption that extends the CAECB assumption to accommodate temporal short-term outcomes. Our proposed assumption states a functional relationship between sequential confounding biases across temporal short-term outcomes, under which we theoretically establish the identification of long-term causal effects. Based on the identification result, we develop an estimator and conduct a theoretical analysis of its asymptotic properties. Extensive experiments validate our theoretical results and demonstrate the effectiveness of the proposed method.
1 Introduction
Long-term causal inference is an important but challenging problem across various scientific fields, such as education (Athey et al., 2019), medicine (Fleming et al., 1994), and marketing (Hohnhold et al., 2015). While in many real-world scenarios, long-term observational data are readily available, a major challenge for long-term causal inference is the presence of latent confounding in observational studies. A common way to mitigate this issue is incorporating short-term experimental data, which raises a fundamental question: how can short-term experimental data be leveraged to address latent confounding in observational data for long-term causal inference, as shown in Figure 1?
Figure 1: Causal graphs of experimental data and observational data with $\mathbf{X}$ being covariates, $\mathbf{U}$ being latent confounders, $A$ being treatment, $\mathbf{S}$ being short-term outcome, and $Y$ being long-term outcome. Fig. 1(a) represents the causal graph of the short-term experimental data, where treatment $A$ is not affected by latent confounders $\mathbf{U}$ and the long-term outcome $Y$ is unobserved. Fig. 1(b) represents the causal graph of the long-term observational data, where the latent confounders $\mathbf{U}$ affect treatments $A$ and outcomes $\mathbf{S},Y$ and the long-term outcome $Y$ can be observed. Fig. 1(c) represents the overall graph with expanding temporal short-term outcomes.
Existing works explore various methods to mitigate latent confounding by utilizing observational and experimental data based on different assumptions. One widely used assumption is the Latent Unconfoundedness assumption (Athey et al., 2020; Chen & Ritzwoller, 2023; Yang et al., 2024), which posits the short-term potential outcomes mediate the long-term potential outcome in the observational data, i.e., $Y(a){\perp\!\!\!\perp}A|\mathbf{S}(a),\mathbf{X},G=O$. These methods are effective when short-term outcomes $\mathbf{S}$ contain substantial information about the latent confounder $\mathbf{U}$, such as in studies on the lifetime effects of youth employment and training programs in the United States (Aizer et al., 2024). However, this assumption essentially restricts that the latent confounders can not affect the long-term outcome, i.e., ruling out the causal edge $\mathbf{U}\rightarrow Y$. To address this limitation, a follow-up work (Ghassami et al., 2022) proposes a novel assumption called Conditional Additive Equi-Confounding Bias (CAECB), allowing for the causal link $\mathbf{U}\rightarrow Y$. The CAECB assumption states that the confounding bias in the short-term outcome is equal to that in the long-term outcome, enabling the identification of long-term causal effects. This assumption may be more reasonable and more aligned with practical settings where confounding biases share a similar manner over short- and long-term outcomes conditional on covariates $\mathbf{X}$, such as in research studying the impact of school finance reforms on student outcomes (Jackson, 2016).
Following the work of Ghassami et al. (2022), we focus on the setting where the causal link $\mathbf{U}\rightarrow Y$ exists. Beyond this, we further consider temporal short-term outcomes, as shown in Figure 1(c), a more common scenario where the method under the CAECB assumption (Ghassami et al., 2022) is not applicable. In many real-world applications, short-term outcomes exhibit temporal dependencies, and capturing these relationships is essential for inferring long-term causal effects. For example, in evaluating the long-term (e.g. year-long) effectiveness of the medication, patients undergo regular follow-up visits (e.g., weekly or monthly). During these visits, temporal short-term health indicators are recorded as short-term outcomes. These sequential measurements capture the progression of patients’ conditions over time, providing valuable information on the long-term outcome of interest. Consequently, the absence of temporal considerations in existing methods limits their effectiveness and constrains the potential of long-term causal inference in practice.
In this paper, we introduce a novel assumption called Functional CAECB (FCAECB), which extends the existing CAECB assumption to capture the temporal dependencies among temporal short-term outcomes. Roughly speaking, the proposed FCAECB assumption posits a functional relationship between sequential latent confounding biases. Under this assumption, we establish the theoretical identification of long-term causal effects. Correspondingly, we devise an algorithm for the long-term effect estimation. Additionally, we analyze the convergence rates of our proposed estimator in the asymptotic and finite sample setting within a generic nonparametric regression framework, with the ultimate goal of deepening the understanding of how sequential short-term confounding biases contribute to inferring long-term effects. Overall, our contribution can be summarized as follows:
  • Novel Assumption for Long-Term Causal Inference: We study the problem of long-term causal inference in the presence of temporal short-term outcomes. We propose a novel assumption named FCAECB for identifying long-term causal effects, which enables capturing the time-dependent relationships between sequential latent confounding biases. Note that the existing CAECB assumption can be seen as our special case.
  • Estimator for Heterogeneous Long-Term Effects: We devise an estimator for estimating heterogeneous long-term effects under the proposed FCAECB assumption, which can be implemented using any machine learning regression method. Theoretically, we analyze the asymptotic properties of our estimator.
  • Empirical Validation: We conduct extensive experiments to validate our theoretical findings and demonstrate the effectiveness of the proposed estimator.
2 Related Work
Long-term causal inference For years, researchers have investigated which short-term outcomes can reliably predict long-term causal effects. Various criteria have been proposed for identifying valid surrogates, including the Prentice criteria (Prentice, 1989), principal surrogacy (Frangakis & Rubin, 2002), strong surrogate criteria (Lauritzen et al., 2004), causal effect predictiveness (Gilbert & Hudgens, 2008), and consistent surrogate and its variations (Chen et al., 2007; Ju & Geng, 2010; Yin et al., 2020). Recently, there has been growing interest in estimating long-term causal effects using surrogates, which is also the focus of this paper. One line of work assumes the unconfoundedness assumption. Under the unconfoundedness assumption, LTEE (Cheng et al., 2021) and Laser (Cai et al., 2024) are based on specifically designed neural networks for long-term causal inference. EETE (Kallus & Mao, 2020) explores the data efficiency from the surrogate in several settings and proposes an efficient estimator for treatment effect. ORL (Tran et al., 2023) introduces a doubly robust estimator for average treatment effects using only short-term experiments, additionally assuming stationarity conditions between short-term and long-term outcomes. Singh & Sun (2024) proposes a kernel ridge regression-based estimator for long-term effect under continuous treatment. Additionally, Wu et al. (2024) develop a policy learning method for balancing short-term and long-term rewards. Our work is different from them. We do not assume the unconfoundedness assumption, and we use the data combination technique to solve the problem of unobserved confounders. Another line of research, which avoids the unconfoundedness assumption, tackles the latent confounding problem by combining experimental and observational data. This setting is first introduced by the method proposed by Athey et al. (2019), which, under surrogacy assumption, constructs the so-called Surrogate Index (SInd) as the substitutions for long-term outcomes in the experimental data for effect identification. As follow-up work, Athey et al. (2020) introduces the latent unconfoundedness assumption, which assumes that short-term potential outcomes can mediate the long-term potential outcomes, thereby enabling long-term causal effect identification. Under this assumption, Yang et al. (2024); Chen & Ritzwoller (2023) propose several estimators for effect estimation. The alternative feasible assumptions (Ghassami et al., 2022) are proposed to replace the latent unconfoundedness assumption, e.g., the Conditional Additive Equi-Confounding Bias (CAECB) assumption. Based on proximal methods (Tchetgen Tchetgen et al., 2024), Imbens et al. (2024) proposes considering the short-term outcomes as proxies of latent confounders, thereby achieving effect identification. However, these studies primarily concentrate on average treatment effects, whereas our focus in this paper is on heterogeneous effects. Among the existing literature, the most closely related work is Ghassami et al. (2022)’s work. Our work can be viewed as a significant extension of theirs, as we consider a more practical scenario where short-term outcomes exhibit temporal dependencies, and theoretically, their CAECB assumption is a special case of the assumption we propose.
Modeling Latent Confounding Bias An effective way to solve the latent confounding problem in the data combination setting is to model latent confounding bias. Kallus et al. (2018) proposes modeling confounding bias under a linearity assumption. Hatt et al. (2022) introduce to model the nonlinear confounding bias using the representation learning technique. Wu & Yang (2022) propose the integrative R-learner via a regularization for the conditional effects and confounding bias with the Neyman orthogonality. Zhou et al. (2025) propose a two-stage representation learning strategy to model such a confounding bias. Different from these works, we focus on the long-term causal inference setting, and rather than focus on how to model the confounding bias, we concentrate more on the relationship between sequential confounding biases.
3 Preliminaries
3.1 Notations, Problem Definition, Assumptions
Let $A\in\{0,1\}$ be the treatment variable, $\mathbf{X}\in{\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ be the observed covariates where $d$ is the dimension of $\mathbf{X}$, $\mathbf{U}\in{\mathcal{U}}\subseteq{\mathbb{R}}^{d_{U}}$ be the latent confounders where $d_{U}$ is the dimension of $\mathbf{U}$. Let $\mathbf{S}=[S_{1},S_{2},\dots,S_{T}]$ be the short-term outcome variable where $S_{t}\in{\mathcal{S}}\subseteq{\mathbb{R}}$ is the short-term outcome measured at timestep $t$, and $Y=S_{T+\mu}\in{\mathcal{R}}\subseteq{\mathbb{R}}$ be the long-term outcome. Further, we leverage the potential outcome framework proposed by Rubin (1978). We denote $\mathbf{S}(a)$ as the potential short-term outcome, $S_{t}(a)$ as the potential short-term outcome at timestep $t$, and $Y(a)$ as the potential long-term outcome. Following existing work on long-term inference (Athey et al., 2020; Ghassami et al., 2022; Imbens et al., 2024; Chen & Ritzwoller, 2023), we denote $G\in\{E,O\}$ as the indicator of data group, where $G=E$ indicates experimental data and $G=O$ indicates observational data. Let lowercase letters (e.g., $a,\mathbf{x},\mathbf{u},\mathbf{s},y,s(a),y(a)$) denote the value of the previously described random variables. Let the superscript $(i)$ denote a specific unit, e.g., $\mathbf{x}^{(i)}$ is the covariate value of unit $i$. Then, the experimental data and the observational data are denoted as $\mathbb{D}_{e}=\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},g^{(i)}=E\}_{i=1}^{n_{e}}$ and $\mathbb{D}_{o}=\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},y^{(i)},G^{(i)}=O\}_{i=n_{e}+1}^{n_{e}+n_{o}}$, where $n_{e}$ and $n_{o}$ are the size of experimental data and the observational data respectively.
For ease of convenience, we denote the following nuisance functions and confounding bias:
$\mu_{S_{t}}^{E}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}|A,\mathbf{X},G=E],$ (1)
$\mu_{S_{t}}^{O}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}|A,\mathbf{X},G=O],$
$\mu_{Y}^{E}(A,\mathbf{X})={\mathbb{E}}[Y|A,\mathbf{X},G=E],$
$\mu_{Y}^{O}(A,\mathbf{X})={\mathbb{E}}[Y|A,\mathbf{X},G=O],$
$\omega_{t}(\mathbf{X})=\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X}),$
where $\omega_{t}(\mathbf{X})$ is known as the confounding bias, the discrepancy between the conditional mean outcome differences derived from the experimental data and the observational data (see Sections 3.2 and 4 for more details on how it serves to the identification of long-term effects).
Moreover, we denote stochastic boundedness with $O_{p}$ and convergence in probability with $o_{p}$. We denote $X_{1}{\perp\!\!\!\perp}X_{2}$ as the independence between $X_{1}$ and $X_{2}$. We use $a_{n}\asymp b_{n}$ to denote both $a_{n}/b_{n}$ and $b_{n}/a_{n}$ are bounded. We use $a_{n}\lesssim b_{n}$ to denote both $a_{n}\leq Cb_{n}$ for some constant $C>0$.
Task: Given a short-term experimental dataset $\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},g^{(i)}=E\}_{i=1}^{n_{e}}$ and a long-term observational dataset $\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},y^{(i)},g^{(i)}=O\}_{i=n_{e}+1}^{n_{e}+n_{o}}$, the goal is to identify the following causal estimand of interest: the Heterogeneous Long-term Causal Effects (HLCE), i.e.,
$\tau(\mathbf{x})={\mathbb{E}}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}].$ (2)
However, HLCE $\tau(\mathbf{x})$ is not identifiable from the experimental data alone, since $Y$ is missing in that dataset. Also it is not identifiable from the observational data alone, since the observational data suffers from the latent confounding problem. Furthermore, the information regarding the causal effects in the experimental data is not necessarily relevant to that in the observational data without further assumptions. To ensure the identification, we first make the following assumptions that are commonly assumed in long-term causal inference:
Assumption 3.1 (Consistency).
If a unit is assigned treatment, we observe its associated potential outcome. Formally, if $A=a$, then $Y=Y(a)$, and $\mathbf{S}=\mathbf{S}(a)$.
Assumption 3.2 (Positivity).
The treatment assignment is non-deterministic. Formally, $\forall a,\mathbf{x}$, we have $10<P(A=a|\mathbf{X}=\mathbf{x})<$ 1, and $0<P(G=O|A=a,\mathbf{X}=\mathbf{x})<1$.
Assumption 3.3 (Weak internal validity of observational data).
Latent confounders exist in observational data. Formally, afor-$\forall a$, we have $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},\mathbf{U},G=O$ and $A\not{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},G=O$.
Assumption 3.4 (Internal validity of experimental data).
There are no latent confounders in experimental data. Formally, afor-$\forall a$, we have $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},G=E$.
Assumption 3.5 (External validity of experimental data).
The distribution of the potential outcomes is invariant to whether the data belongs to the experimental or observational data. Formally, afor-$\forall a$, we have $G{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X}$.
Assumptions 3.1 and 3.2 are standard assumptions in causal inference (Rubin, 1974; Imbens, 2000). Assumptions 3.3, 3.4 and 3.5 are mild and widely assumed in data combination settings (Shi et al., 2023; Imbens et al., 2024; Athey et al., 2019, 2020; Hu et al., ). Specifically, Assumption 3.3 allows the existence of latent confounders in observational data, thus it is much weaker than the traditional unconfoundedness assumption. Assumption 3.4 is reasonable and can hold since the treatment assignment mechanism is under control in the experiments. Assumption 3.5 connects the potential outcome distributions between observational and experimental data.
Assumptions 3.1, 3.2, 3.3, 3.4 and 3.5 are still not sufficient to identify the causal estimand of interest. The root cause is that, even though the assumptions above link the experimental and observational data, the (long-term) latent confounding problem remains unsolved. In the following section, we first review a method proposed by Ghassami et al. (2022), which poses an extra assumption, called Conditional Additive Equi-Confounding Bias (CAECB) assumption, to achieve the identification of HLCE. Then, we propose our approach that generalizes the method of Ghassami et al. (2022) to allow temporal short-term outcomes.
(a) Schematic representation of CAECB assumption (Ghassami et al., 2022).
(b) Schematic representation of our proposed FCAECB assumption.
Figure 2: Schematic representations of CAECB assumption proposed by Ghassami et al. (2022) and our FCAECB assumption. As shown in Figure 2(a), the CAECB assumption requires that the confounding bias in the short-term outcome is equal to that in the long-term outcome. As shown in Figure 2(b), the FCAECB assumption relaxes this constraint by allowing for temporal short-term outcomes and only requiring that confounding biases across different time steps follow a specific pattern rather than remaining equal.
3.2 CAECB Assumption (Ghassami et al., 2022)
Ghassami et al. (2022) introduced a method for combining experimental and observational data to identify the HLCE under an extra CAECB assumption. The method under CAECB assumption is only applicable to scenarios with a one-dimensional short-term outcome, which we denote as $S$ (and its corresponding potential outcome $S(a)$) with slightly abusing notation.
To begin with, we first restate the CAECB assumption as follows:
Assumption 3.6 (Conditional Additive Equi-Confounding Bias, CAECB (Ghassami et al., 2022)).
The difference in conditional expected values of short-term potential outcomes across treated and control groups is the same as that of the long-term potential outcome variable. Formally, afor-$\forall a$, we have
[…]
Theorem 3.7.
Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6 hold, then $\tau(\mathbf{x})$ can be identified as follows:
$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$ (4)
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+\omega(\mathbf{x}),$
where $\omega(\mathbf{x})=\mu_{S}^{E}(1,\mathbf{x})-\mu_{S}^{E}(0,\mathbf{x})+\mu_{S}^{O}(0,\mathbf{x})-\mu_{S}^{O}(1,\mathbf{x})$ is the short-term confounding bias.
Proof can be found in Appendix A. A similar identification result in terms of long-term average causal effects has been shown by Ghassami et al. (2022). Here we provide the identification result of HLCE $\tau(\mathbf{x})$ in Theorem 3.7.
Remark 3.8.
We illustrate Assumption 3.6 in Figure 2(a). Assumption 3.6 can be seen as a generalization of the parallel assumption in the difference-in-differences (DiD) framework (Ashenfelter & Card, 1984; Angrist & Pischke, 2009), Note that Assumption 3.6 essentially implies that the short-term confounding bias is the same as the long-term one. In this way, $\omega(\mathbf{x})$ can be interpreted as the long-term confounding bias, resulting in the identification result in Eq. (4).
Considering Eq. (3) in Assumption 3.6, $S$ and $Y$ should be on the same scale, which restricts the practical application of the method under this assumption. In the next section, we extend this assumption to allow for temporal short-term outcomes, which significantly improve the practical utility.
4 Long-term Identification under Functional CAECB Assumption
In this section, we introduce a novel assumption considering the temporal information between temporal short-term outcomes, enabling the identification of long-term causal effects.
Specifically, we formalize our proposed assumption:
Assumption 4.1 (Functional Conditional Additive Equi-Confounding Bias, FCAECB).
The difference in conditional expected values of short-term potential outcomes across treated and control groups between two timesteps follows the learnable function forms $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$. Formally, afor-$\forall a$, we have
$b_{t+1}(a,\mathbf{X})=f(\mathbf{X})b_{t}(a,\mathbf{X}),$ (5)
where $b_{t}(a,\mathbf{X})={\mathbb{E}}[S_{t}(a)|\mathbf{X},A=0,G=O]-{\mathbb{E}}[S_{t}(a)|\mathbf{X},A=1,G=O]$.
Remark 4.2.
We illustrate our proposed FCAECB assumption in Figure 2(b). We relax the CAECB assumption (Ghassami et al., 2022) to allow for the temporal short-term outcomes, instead of restricting that the short-term outcome should be the same scale as the long-term outcome. Additionally, the existing CAECB assumption can be seen as our special case when $f(\mathbf{X})$ in the FCEACB assumption satisfies $f(\mathbf{X})=1$.
To provide a better understanding of our Assumption 4.1, we provide the insight in term of the functional form $f(\mathbf{X})$ in the following proposition.
Proposition 4.3.
Under Assumption 4.1, $\forall t$ the confounding biases between times $t$ and $t+1$ follow
$\omega_{t+1}(\mathbf{X})=f(\mathbf{X})\omega_{t}(\mathbf{X}),$ (6)
where $\omega_{t}(\mathbf{X})$ is the confounding bias at time step $t$, defined as $\omega_{t}(\mathbf{X})=\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X})$.
The proof is given in Appendix B.
Remark 4.4.
As stated in Proposition 4.3, Assumption 4.1 essentially states the confounding biases $\omega_{t}(\mathbf{X})$ between adjacent time steps follow the functional form $f(\mathbf{X})$. Proposition 4.3 also illustrates the way of how to learn the function $f(\mathbf{X})$ using the observed variables, unlike the definition in Eq. (5) defining $f(\mathbf{X})$ using the potential outcomes (See Section 5 for the estimation based on Proposition 4.3).
The key to identifying long-term causal effect under Assumption 4.1 is using the temporal information in $f(\mathbf{X})$ to extrapolate the long-term confounding bias $\omega_{T+\mu}(\mathbf{X})$. Using Eq. (6), the long-term confounding bias $\omega_{T+\mu}(\mathbf{X})$ can be expressed as $\omega_{T+\mu}(\mathbf{X})=f^{\mu}(\mathbf{X})\omega_{T}(\mathbf{X})$, which results in the following theorem of long-term effect identification.
Theorem 4.5.
Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 4.1 hold, then the heterogeneous long-term effects $\tau(\mathbf{x})$ can be identified as follows:
$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$ (7)
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x}),$
where $\omega_{T}(\mathbf{x})=\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})$ is the short-term confounding bias at time step $T$.
The proof is given in Appendix C. Theorem 4.5 provides the identification result of the heterogeneous long-term effects $\tau(\mathbf{X})$. The identification result consists of two parts: 1. long-term outcome differences in observational data $\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})$, and 2. long-term confounding bias $f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x})$. The long-term confounding bias is identified by the extrapolated result using short-term confounding bias $\omega_{T}(\mathbf{x})$ under Assumption 4.1. The identification result also illustrates a way to estimate the long-term effects $\tau(\mathbf{x})$ via modeling sequential latent confounding, which inspires our estimator as shown in Section 5.
5 Long-term Causal Effect Estimation
In this section, we first introduce our estimator $\hat{\tau}(\mathbf{X})$ for heterogeneous long-term effects under our proposed Assumption 4.1, and provide the corresponding theoretical analysis of the proposed estimator.
5.1 Estimator
Our estimator $\hat{\tau}(\mathbf{X})$ directly follows the identification result in Theorem 4.5. As shown in Eq. (7), the estimators consist of three nuisance components, outcome mean difference between treated and control group in observational data $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})$, confounding bias $\omega_{T}(\mathbf{x})$, and the function between confounding bias $f(\mathbf{x})$. The first and the second terms can be directly estimated by fitting nuisance functions, and the third term $f(\mathbf{x})$ should be estimated based on the fitted confounding biases. Accordingly, we design our heterogeneous long-term effect estimator within a two-stage regression framework, which are model-agnostic algorithms that decompose the estimation task into multiple sub-problems, each solvable using any supervised learning/regression methods.
Specifically, our proposed estimator follows:
  • S0.
    (Optional) Selecting subsets of short-term outcomes $\mathbf{S}$, yielding appropriate $T$ and $\mu$;
  • S1.
    Fitting the following nuisance functions: $\hat{\mu}_{Y}^{O}(a,\mathbf{x})$, $\hat{\mu}_{S_{t}}^{E}(a,\mathbf{x})$, and $\hat{\mu}_{S_{t}}^{O}(a,\mathbf{x})$ for all $a$ and $t$;
  • S2.
    Constructing the confounding bias $\hat{\omega}_{t}(\mathbf{x})=\hat{\mu}_{S_{t}}^{E}(1,\mathbf{x})-\hat{\mu}_{S_{t}}^{E}(0,\mathbf{x})+\hat{\mu}_{S_{t}}^{O}(0,\mathbf{x})-\hat{\mu}_{S_{t}}^{O}(1,\mathbf{x})$ for all $t$, and fitting the function $\hat{f}(\mathbf{x})$ by minimizing
    $arg\!\min_{f\in{\mathcal{F}}}\Sigma_{t=1}^{T-1}\Sigma_{i=1}^{n}\left(\hat{\omega}_{t+1}(\mathbf{x}_{i})-f(\mathbf{x}_{i})\hat{\omega}_{t}(\mathbf{x}_{i})\right)^{2};$ (8)
  • S3.
    Constructing final HLCE estimator as
    $\hat{\tau}(\mathbf{x})=\hat{\mu}_{Y}^{O}(1,\mathbf{x})-\hat{\mu}_{Y}^{O}(0,\mathbf{x})+\hat{f}^{\mu}(\mathbf{x})$ (9)
    ×$\times\left(\hat{\mu}_{S_{T}}^{E}(1,\mathbf{x})-\hat{\mu}_{S_{T}}^{E}(0,\mathbf{x})+\hat{\mu}_{S_{T}}^{O}(0,\mathbf{x})-\hat{\mu}_{S_{T}}^{O}(1,\mathbf{x})\right).$
Note that, in addition to the fitting/constructing steps S1-S3, we also introduce an optional S0 to select subsets of short-term outcomes $\mathbf{S}$. This step is motivated by the identification result in Eq. (7) that allows for different choices of $\mu$ and $T$. For example, suppose we can observe $6$-step short-term outcome $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ and the long-term outcome of interest is $Y=S_{9}$. Then we have multiple choices of $T$ and $\mu$, e.g., using all short-term outcomes $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ with $T=6$ and $\mu=3$, or using $\mathbf{S}=[S_{1},S_{3},S_{5}]$ with $T=3$ and $\mu=2$, or using $\mathbf{S}=[S_{1},S_{5}]$ with $T=2$ and $\mu=1$. We will discuss how to choose $T$ and $\mu$ in the next section.
5.2 Convergence Rate Analyses
In this paper, we assume the smoothness assumption of the estimated functions, where $s$-smooth functions are contained in the Hölder ball $\mathcal{H}_{d}(s)$, estimable with the minimax rate (Stone, 1980) of $n^{\frac{1}{2+d/s}}$ where $d$ is the dimension of $\mathcal{X}$. Formally, we provide the following definition.
Definition 5.1 (Hölder ball).
The Hölder ball $\mathcal{H}_{d}(s)$ is the set of $s$-smooth functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ supported on $\mathcal{X}\subseteq\mathbb{R}^{d}$ that are $\lfloor s\rfloor$-times continuously differentiable with their multivariate partial derivatives up to order $\lfloor s\rfloor$ bounded, and for which
$|\frac{\partial^{m}f}{\partial^{m_{1}}\cdots\partial^{m_{d}}}(x)-\frac{\partial^{m}f}{\partial^{m_{1}}\cdots\partial^{m_{d}}}(x^{\prime})|\lesssim\|x-x^{\prime}\|^{s-\lfloor s\rfloor}_{2},$
$\forall x,x^{\prime}$ and $m=(m1,\cdots,m_{d})$ such that $\Sigma_{j=1}^{d}m_{j}=\lfloor s\rfloor$.
Assumption 5.2 (Smoothness Assumption).
We assume that the nuisance functions and the mapping $f$ defined in Assumption 4.1 satisfy: (1) $\mu_{S_{t}}^{E}$, $\mu_{S_{t}}^{O}$, and $\mu_{Y}^{O}$ are $\alpha$-smooth, $\beta$-smooth, and $\gamma$-smooth, respectively, and all are estimable at Stone (1980)’s minimax rate of $n^{\frac{-p}{2p+d}}$ for a $p$-smooth function; (2) $f(x)$ is $\eta$-smooth.
We also assume the boundedness assumption of the nuisance functions.
Assumption 5.3 (Boundness Assumption).
We assume that the nuisance functions $\mu_{S_{t}}^{O}(a,x),\mu_{S_{t}}^{E}(a,x)$ and their estimates are bounded, i.e., $\forall t$, $|\mu_{S_{t}}^{O}(a,x)|<C_{1}$, $|\hat{\mu}_{S_{t}}^{O}(a,x)|<C_{2}$, $|\mu_{S_{t}}^{E}(a,x)|<C_{3}$ and $|\hat{\mu}_{S_{t}}^{E}(a,x)|<C_{4}$ hold for some $C_{1},C_{2},C_{3},C_{4}>0$.
We now state our main theoretical results. To obtain our error bounds of the proposed estimator, we leverage the same sample splitting technique from (Kennedy, 2023), which randomly splits the datasets into two independent sets and applies them to the regressions of the first step and second step respectively. Such a technique is originally used to analyze the convergence rate of the double robust conditional average treatment effect estimation in the traditional setting (Kennedy, 2023) and later is adapted to several other methods (Curth & Van der Schaar, 2021; Frauen & Feuerriegel, ). Different from them, we use such a technique for the sequential latent confounding modeling, which is then adapted for the long-term effect estimation.
To begin with, we first provide the rate of $\hat{f}(\mathbf{x})$ as follows:
Lemma 5.4.
Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, and 5.2 hold, then we have
$\hat{f}(\mathbf{x})-$ $f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}\right.$ (10)
$\left.+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$
The proof is given in Appendix D. Since $\hat{f}(\mathbf{x})$ is a time-series model, its effective sample is $(T-1)n$, thus it might achieve a faster rate if we observed a longer duration of short-term outcomes. Moreover, the rate of $\hat{f}(\mathbf{x})$ consist two part: the oracle rate $(\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}$, and the rate of fitting nuisance functions $(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}$. If the nuisance functions $\mu_{S_{t}}^{E}$ and $\mu_{S_{t}}^{O}$ is smooth enough such that $\min\{\alpha,\beta\}\geq\eta$, then $\hat{f}(\mathbf{x})$ will attain the oracle rate.
Based on Lemma 5.4, we provide the rate of $\hat{\tau}(\mathbf{x})$ in the following theorem.
Theorem 5.5.
Suppose Lemma 5.4 hold, then we have
$\hat{\tau}(\mathbf{x})-\tau(\mathbf{x})$ (11)
$=$ $O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}\right.$
$\left.+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$
Proof can be found in Appendix E. Theorem 5.5 follows directly from Lemma 5.4. The rate of $\hat{\tau}(\mathbf{x})$ consist of several terms: 1. $n^{-\frac{\gamma}{2\gamma+d}}$ represents the rate of $\mu_{Y}^{O}(a,\mathbf{x})$, 2. $n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}$ corresponds the rate of confounding bias $\omega_{T}(\mathbf{x})$, 3. and the remaining term is the rate of $\hat{f}^{\mu}(\mathbf{x})$. Theorem 5.5 suggests that the convergence rate is primarily influenced by the smoothness parameters $\alpha,\beta,\gamma,\eta$ and the temporal parameters $T,\mu$, where $T$ represents the longest observed duration of the short-term outcomes $\mathbf{S}$ and $\mu$ represents the time horizon of the long-term outcome to be estimated. This implies that achieving a more accurate effect estimation requires observing short-term outcomes for as long as possible.
Remark 5.6 (Choosing $\mu$ and $T$).
Note that, the bound in Eq. (11) contains the term $\frac{\mu}{(T-1)^{\frac{\alpha}{2\alpha+d}}}$ (similar for $\beta$ and $\eta$). If we have prior knowledge of the smoothness of nuisance functions, i.e., knowing $\alpha$ ($\beta$ and $\eta$), then we can optimally select $\mu$ and $T$ to obtain a faster rate. For instance, if all estimated functions are sufficiently smooth, i.e., $\alpha,\beta,\eta\rightarrow\infty$, then $\mu$ and $T$ should be chosen to minimize $\frac{\mu}{\sqrt{T-1}}$, since $\frac{\alpha}{2\alpha+d},\frac{\beta}{2\beta+d},\frac{\eta}{2\eta+d}\rightarrow 1/2$. As a concrete example, suppose we can observe $6$-step short-term outcome $\mathbf{S}=[S_{1},S_{2},\dots,S_{6}]$ and the long-term outcome of interest is $Y=S_{9}$. The optimal choice is to use $\mathbf{S}=[S_{1},S_{5}]$ only to estimate $\tau(\mathbf{x})$, since it result in a minimum $\frac{\mu}{\sqrt{T-1}}=1$ where $\mu=1$ and $T=2$. In practice, when no prior knowledge is available, it may be advisable to choose $\mu$ as small as possible, since a larger $\mu$-power in $\hat{f}^{\mu}(\mathbf{x})$ could lead to significant errors.
6 Experiments
In this section, we perform a series of experiments to evaluate the effectiveness and validity of our proposed estimator. We first provide the experimental setup in Section 6.1, and provide the experimental results and corresponding analyses in Section 6.2.
6.1 Experimental Setup
In this section, we describe the data generation process for the synthetic datasets, the implementation details, and the metrics used in the experiments.
Synthetic Datasets The data generation process is partly following (Kallus et al., 2018) such that we can obtain a specific form of confounding bias. Specifically, we first generate the treatments as follows: $A|G=O\sim\text{Bernoulli}(0.6)$ and $A|G=E\sim\text{Bernoulli}(0.4)$. Then we generate the observed $\mathbf{X}$ and the unobserved $\mathbf{U}$ as follows:
$(\mathbf{X},\mathbf{U})|A,G=E\sim\mathcal{N}([\frac{2A-1}{2},0],\begin{bmatrix}1&0\\ 0&1\end{bmatrix})$ (12)
$(\mathbf{X},\mathbf{U})|A,G=O\sim\mathcal{N}([\frac{1-2A}{2},0],\begin{bmatrix}1&A-0.5\\ A-0.5&1\end{bmatrix}).$
Finally, we generate the $T$-step short-term outcomes $\mathbf{S}$ and the long-term outcome $Y=S_{\mu}$ satisfying Assumption 4.1 as follows:
$S_{t}(1)=1+1.1\mathbf{X}+\mathbf{U}+\sum_{k=0}^{t}S_{k}(1)+\epsilon_{S_{t}},$ (13)
$S_{t}(0)=\mathbf{X}+\mathbf{U}+\sum_{k=0}k^{t}S_{k}(0)+\epsilon_{S_{t}},$
where $\epsilon_{S_{t}}$ are Gaussian noises. This will result in non-equal confounding bias in different time steps $t$, i.e., $\omega_{t}(\mathbf{x})=t\times\mathbf{x}$. We perform several control experiments using this data generation process. The default values are: the observational data sample size $n_{o}=4000$, experimental data sample size $n_{e}=2000$, $T=6$, and $\mu=3$. In control experiments, we vary $\mu$ within $\{1,2,3,4,5\}$ and $T$ within $T\in\{4,5,6,7,8\}$. We also fix the ratio of $n_{O}:n_{E}=2:1$, and varying $n_{e}$ within $\{1000,2000,3000,4000,5000,10000\}$ (which corresponds to $n_{o}$ values of $\{2000,4000,6000,8000,10000,20000\}$).
Implementation We fit all nuisance functions using correctly specified regression. We denote our estimator as $\hat{\tau}$. For comparison, we evaluate our method against a correctly specified T-Learner (Künzel et al., 2019), denoted as $\hat{\tau}_{exp}$, which directly regresses unobserved $Y$ in the experimental data. However, this estimator $\hat{\tau}_{exp}$ is infeasible in practice since $Y$ is missing in the experimental data. This baseline serves as an idealized benchmark, and comparing our method against it highlights the effectiveness of our estimator across different control experiments.
Metrics As for heterogeneous effect estimation, we report Precision in the Estimation of Heterogeneous Effect (PEHE) $\varepsilon_{PEHE}=\sqrt{\frac{1}{n}\Sigma_{i=1}^{n}(\tau(\mathbf{x}_{i})-\hat{\tau}(\mathbf{x}_{i}))^{2}}$. As for average long-term causal effect estimation, we report the absolute error $\varepsilon_{ATE}=|\Sigma_{i=1}^{n}\tau(\mathbf{x}_{i})-\Sigma_{i=1}^{n}\hat{\tau}(\mathbf{x}_{i}))|$. For all metrics, we report the mean values and standard errors by $100$ times running.
6.2 Experimental Results
(a) Heterogeneous effect estimation error.
(b) Average effect estimation error.
Figure 3: Results of the experiments in terms of different choice of $\mu$ and $T$.
Optimal Choice of $T$ and $\mu$: We conduct experiments using different choice of $T$ and $\mu$. Specifically, to predict long-term effects, we consider the following subsets of $\mathbf{S}$: 1. $\mathbf{S}=[S_{1},S_{3}]$ with $T=2,\mu=3$; 2. $\mathbf{S}=[S_{1},S_{3},S_{5}]$ with $T=3,\mu=2$; 3. $\mathbf{S}=[S_{1},S_{2},S_{3},S_{4},S_{5}]$ with $T=5,\mu=3$; and 4. $\mathbf{S}=[S_{1},S_{5}]$ with $T=2,\mu=1$. The results, shown in Figure 3, are sorted by the values of $\frac{\mu}{\sqrt{T-1}}$, which $3,1.41,1.34$, and $1$, respectively. As shown in Figure 3, as $\frac{\mu}{\sqrt{T-1}}$ decrease, both the estimation errors $\varepsilon_{PEHE}$ and $\varepsilon_{ATE}$ decrease. This aligns with our theoretical findings in Theorem 5.5, as a smaller $\frac{\mu}{\sqrt{T-1}}$ leads to a faster rate. Interestingly, as shown in Figure 5(b), we found that when the $\frac{\mu}{\sqrt{T-1}}$ is equal to $1$, our estimator $\hat{\tau}_{(}\mathbf{x})$ achieves lower $\varepsilon_{ATE}$ than the idealized estimator $\hat{\tau}$. This may be attributed to the fact that our estimator $\hat{\tau}(\mathbf{x})$ leverages observational data, leading to higher data efficiency.
(a) Heterogeneous effect estimation error with fixed $T$ and varying $\mu$.
(b) Average effect estimation error with fixed $T$ and varying $\mu$.
(c) Heterogeneous effect estimation error with fixed $\mu$ and varying $T$.
(d) Average effect estimation error with fixed $\mu$ and varying $T$.
Figure 4: Results of control experiments with fixed $T$ and varying $\mu$ (Figures 4(a) and 4(b)), and with fixed $\mu$ and varying $T$ (Figures 4(c) and 4(d)).
Vary $T$ and $\mu$: We conduct control experiments with varying $T$ and varying $\mu$ separately. First, Figures 4(a) and 4(b)) show the results for the experiments with fixed $T$ and varying $\mu$. As expected, the estimation errors on both heterogeneous and average effect estimations increase as the long-term horizon $\mu$ grows. This is primarily due to the increasing estimation errors of $\hat{f}^{\mu}(\mathbf{x})$, as established in Lemma 5.4. Second, Figures 4(c) and 4(d) show the results for the experiments with fixed $\mu$ and varying $T$. As the observed duration $T$ increases, the estimation errors decrease rapidly. It is reasonable since the larger $T$ results in a faster rate of $\hat{\tau}(\mathbf{x})$. Both findings further support our Lemma 5.4 and Theorem 5.5 again. Moreover, we observe that, with smaller $\mu$, our estimator closely matches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, showing the effectiveness of our estimator.
(a) Heterogeneous effect estimation error.
(b) Average effect estimation error.
Figure 5: Results of control experiments with varying sample sizes $n_{e}$ and $n_{o}$.
Sample sensitivity: We conduct control experiments by varying the sample sizes of both experimental and observational data. The results are presented in Figure 5. As expected, the estimator’s performance improves as the sample sizes $n_{o}$ and $n_{e}$ increase. Notably, when $n_{o}$ and $n_{e}$ become larger, our estimator $\hat{\tau}(\mathbf{x})$ closely approaches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, which again confirms the effectiveness of our estimator. Furthermore, when the sample size $n_{e}\geq 3000$ ($n_{e}\geq 6000$), the estimation errors become stable, indicating an appropriate sample size threshold.
7 Conclusion
In this paper, we extend the existing CAECB assumption (Ghassami et al., 2022) for long-term causal inference. The original assumption is restricted to settings with a one-dimensional short-term outcome. To address this limitation, we introduce a more general assumption—the functional CAECB assumption—which accommodates temporal short-term outcomes. We theoretically establish the identification result of the heterogeneous long-term causal effect under our assumption and propose a corresponding estimator by modeling sequential latent confounding. Additionally, we provide a comprehensive theoretical analysis of the estimator. Experiments confirm both the validity of our theoretical results and the effectiveness of the proposed estimator.
Impact Statement
This paper presents work whose goal is to advance the field of Long-term causal inference in the presence of latent confounders. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proof of Theorem 3.7
We first restate Theorem 3.7 as follows:
Theorem A.1.
Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6 hold, then $\tau(\mathbf{x})$ can be identified as follows:
$\tau(\mathbf{x})$ (14)
$=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+\mu_{S}^{E}(1,\mathbf{x})-\mu_{S}^{E}(0,\mathbf{x})+\mu_{S}^{O}(0,\mathbf{x})-\mu_{S}^{O}(1,\mathbf{x}).$
Proof.
$\tau(\mathbf{X})$ (15)
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]\}\times p(A=0|\mathbf{X},G=O)$
+$+\{\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]\}\times p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}\left[S(1)|\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(1)|\mathbf{X},G=O,A=1\right]\}\times p(A=0|\mathbf{X},G=O)$
+$+\{\mathbb{E}\left[S(0)|\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(0)|\mathbf{X},G=O,A=1\right]\}\times p(A=1|\mathbf{X},G=O),$
where the first equality is based on Assumption 3.5 and the last equality is based on Assumption 3.6. Similarly, for short-term conditional causal effects, we have:
$\mathbb{E}[S(1)|\mathbf{X},G=O]-\mathbb{E}[S(0)|\mathbf{X},G=O]$ (16)
$=$ $\mathbb{E}[S|\mathbf{X},G=O,A=1]-\mathbb{E}[S|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S(1)|\mathbf{X},G=O,A=1]\}\times p(A=0|\mathbf{X},G=O)$
+$+\{\mathbb{E}[S(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S(0)|\mathbf{X},G=O,A=1]\}\times p(A=1|\mathbf{X},G=O)$
Then, combining Eq. (15) and (16), we have
$\tau(\mathbf{X})$ (17)
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S(1)|\mathbf{X},G=O]-\mathbb{E}[S(0)|\mathbf{X},G=O]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S(1)|\mathbf{X},G=E]-\mathbb{E}[S(0)|\mathbf{X},G=E]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S|\mathbf{X},G=E,A=1]-\mathbb{E}[S|\mathbf{X},G=E,A=0]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
$=$ $\mu_{Y}^{O}(1,\mathbf{X})-\mu_{Y}^{O}(0,\mathbf{X})+\mu_{S}^{E}(1,\mathbf{X})-\mu_{S}^{E}(0,\mathbf{X})+\mu_{S}^{O}(0,\mathbf{X})-\mu_{S}^{O}(1,\mathbf{X}),$
where the second equality is based on Assumption 3.5 and the last equality is based on Assumption 3.4. This finishes our proofs. ∎
Appendix B Proof of Proposition 4.3
We first restate the proposition as follow:
Proposition B.1.
Under Assumption 4.1, $\forall t$ the confounding biases between times $t$ and $t+1$ follow
$\omega_{t+1}(\mathbf{X})=f(\mathbf{X})\omega_{t}(\mathbf{X}),$ (18)
where $\omega_{t}(\mathbf{X})$ is the confounding bias at time step $t$, defined as $\omega_{t}(\mathbf{X})=\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X})$.
Proof.
We start from the long-term causal effects at time step $t$ in experimental data $G=E$:
$\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]$ (19)
=$\overset{(a)}{=}$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O]$
$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
$+b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
where the equality $(a)$ is based on Assumption 3.5. By rewriting the last equality above, we obtain:
$\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]$ (20)
$=$ $\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
$+b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
$\iff$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]+\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
$=$ $b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
$\iff$ $\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X})$
$=$ $b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
$\overset{(a)}{\iff}$ $\omega_{t}(\mathbf{X})=b_{t}(1,\mathbf{X}))p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
where $(a)$ is based on the definition of $\omega_{t}(\mathbf{X})$. Then, similarly for time step $t+1$, we have
$\omega_{t+1}(\mathbf{X})$ $=b_{t+1}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t+1}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$ (21)
=$\overset{(a)}{=}f(\mathbf{X})\left(b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)\right)$
=$\overset{(b)}{=}f(\mathbf{X})\omega_{t}(\mathbf{X}),$
where $(a)$ is based on Assumption 4.1, i.e., $b_{t+1}(a,\mathbf{X})=f(\mathbf{X})b_{t}(a,\mathbf{X})$, and $(b)$ is based on Eq. (20) and finishes our proof. ∎
Appendix C Proof of Theorem 4.5
We first restate Theorem 4.5 as follows:
Theorem C.1.
Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 4.1 hold, then $\tau(\mathbf{x})$ can be identified as follows:
$\tau(\mathbf{x})$ (22)
$=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{X})\left(\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})\right).$
Proof.
$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$ (23)
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+b_{T+\mu}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{T+\mu}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)$
where the first equality is based on Assumption 3.5, i.e., $G{\perp\!\!\!\perp}Y(a)|\mathbf{X}$, and last equality is based on Assumption 4.1.
Further based on Assumption 4.1, we have $b_{T+\mu}(a,\mathbf{X})=f^{\mu}(\mathbf{X})b_{T}(a,\mathbf{X})$. Then, we rewrite the equality above as:
$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$ (24)
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+f^{\mu}(\mathbf{X})b_{T}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-f^{\mu}(\mathbf{X})b_{T}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)$
Similarly, for short-term ITE at last time step $T$ we have:
$\mathbb{E}[S_{T}(1)|\mathbf{X}]-\mathbb{E}[S_{T}(0)|\mathbf{X}]$ (25)
$=$ $\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S_{T}(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[S_{T}(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
$+b_{T}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{T}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)$
Then, combining Eq. 24 and 25, we have
$\tau(\mathbf{X})=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(1)|\mathbf{X}]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(0)|\mathbf{X}]$ (26)
$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=E,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=E,A=0]$
$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{X})\left(\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})\right)$
where the second equality is based on Assumption 3.5. This finishes our proof. ∎
Appendix D Proof of Lemma 5.4
We first state the following lemma that is used in the proof of Lemma 5.4.
Lemma D.1.
suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1 hold, then $\forall t$,
$\hat{\omega}_{t}(\mathbf{x})-\omega_{t}(\mathbf{x})=O_{p}\left(r_{\mu_{S_{t}}^{E}}(n)+r_{\mu_{S_{t}}^{O}}(n)\right).$ (27)
where $r_{\circ}(n)$ denotes the risk of nuisance function $\circ$, e.g., $r_{\mu_{S_{t}}^{E}}(n)$ correspondingly to $\mu_{S_{t}}^{E}$, and further under Assumption 5.2, we have
$\hat{\omega}_{t}(\mathbf{x})-\omega_{t}(\mathbf{x})=O_{p}\left(n^{\frac{-\alpha}{2\alpha+d}}+n^{\frac{-\beta}{2\beta+d}}\right).$ (28)
Proof.
The lemma is immediately proved by the form of $\omega_{t}(\mathbf{x})$ as $\omega_{t}(\mathbf{x})=\mu_{S_{t}}^{E}(1,\mathbf{x})-\mu_{S_{t}}^{E}(0,\mathbf{x})+\mu_{S_{t}}^{O}(0,\mathbf{x})-\mu_{S_{t}}^{O}(1,\mathbf{x})$. ∎
We now restate our Lemma 5.4 as follows and formally prove that.
Lemma D.2.
Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, 5.2 and 5.3 hold, then we have
$\hat{f}(\mathbf{x})-f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$ (29)
Proof.
We apply Proposition 1 of Kennedy et al. (Kennedy, 2023), yielding that
$\hat{f}(\mathbf{x})-f(\mathbf{x})=$ $(\hat{f}(\mathbf{x})-\tilde{f}(\mathbf{x}))+(\tilde{f}(\mathbf{x})-f(\mathbf{x}))$ (30)
$=$ $(\hat{f}(\mathbf{x})-\tilde{f}(\mathbf{x}))+O_{p}(R_{n}^{*}(\mathbf{x}))$
$=$ $\hat{\mathbb{E}}_{n}[\hat{r}(\mathbf{X})|\mathbf{X}=\mathbf{x}]+o_{p}(R_{n}^{*}(\mathbf{x}))+O_{p}(R_{n}^{*}(\mathbf{x}))$
where $\hat{r}(\mathbf{x})=\mathbb{E}[\hat{f}(\mathbf{X})|\mathbf{X}=\mathbf{x}]-f(\mathbf{x})$, and $\mathcal{R}_{n}^{*}(x)$ is the oracle risk of second-stage regression and further under Assumption 5.2, we know $f$ is $\eta$-smooth, thus $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$. Here, $f$ is a time-series model ($1$-order autoregressive), optimized with $T-1$-length steps and $n$ samples, as shown in Eq. (8). According to Proposition 4.3, $\forall t$, $f(\mathbf{x})=\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}$. We first prove the rate of $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}$ for a fixed $t$ as follows.
$(\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}-\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})})^{2}$ (31)
$=$ $\left(\frac{\hat{\omega}_{t+1}(\mathbf{x})\omega_{t}(\mathbf{x})-\omega_{t+1}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})}{\omega_{t}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})}\right)^{2}$
$=$ $\left(\frac{\hat{\omega}_{t+1}(\mathbf{x})\omega_{t}(\mathbf{x})-\omega_{t+1}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})+\omega_{t+1}(\mathbf{x})\omega_{t}(\mathbf{x})-\omega_{t+1}(\mathbf{x})\omega_{t}(\mathbf{x})}{\omega_{t}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})}\right)^{2}$
$=$ $\left(\frac{(\hat{\omega}_{t+1}(\mathbf{x})-\omega_{t+1}(\mathbf{x}))\omega_{t}(\mathbf{x})+\omega_{t+1}(\mathbf{x})(\omega_{t}(\mathbf{x})-\hat{\omega}_{t}(\mathbf{x}))}{\omega_{t}(\mathbf{x})\hat{\omega}_{t}(\mathbf{x})}\right)^{2}$
$\overset{(a)}{\leq}$ $\frac{2(\hat{\omega}_{t+1}(\mathbf{x})-\omega_{t+1}(\mathbf{x}))^{2}\omega_{t}^{2}(\mathbf{x})+\omega_{t+1}^{2}(\mathbf{x})(\omega_{t}(\mathbf{x})-\hat{\omega}_{t}(\mathbf{x}))^{2}}{\omega_{t}^{2}(\mathbf{x})\hat{\omega}_{t}^{2}(\mathbf{x})}$
$=$ $\frac{2}{\hat{\omega}_{t}^{2}(\mathbf{x})}(\hat{\omega}_{t+1}(\mathbf{x})-\omega_{t+1}(\mathbf{x}))^{2}+\frac{2\omega_{t+1}^{2}(\mathbf{x})}{\omega_{t}^{2}(\mathbf{x})\hat{\omega}_{t}^{2}(\mathbf{x})}(\omega_{t}(\mathbf{x})-\hat{\omega}_{t}(\mathbf{x}))^{2}$
$\overset{(b)}{\asymp}$ $(\hat{\omega}_{t+1}(\mathbf{x})-\omega_{t+1}(\mathbf{x}))^{2}+(\omega_{t}(\mathbf{x})-\hat{\omega}_{t}(\mathbf{x}))^{2};$
where the inequality $(a)$ is based on $(a+b)^{2}\leq 2(a^{2}+b^{2})$, and $(b)$ is based on Assumption 5.3. Under Assumption 5.3, $\mu_{S_{t}}^{O}$, $\mu_{S_{t}}^{E}$ and their estimates are all bounded, and thus $\hat{\omega}_{t}(\mathbf{x})$ and its estimates are also bounded, thus $(b)$ holds. Then, applying Lemma D.1, we know for a fixed time step $t$, $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}-\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}=O_{p}(n^{\frac{-\alpha}{2\alpha+d}}+n^{\frac{-\beta}{2\beta+d}})$. Then, for $\hat{r}(\mathbf{x})$, we have the effective sample is of size $(T-1)n$, and then combining with Eq. (30) and $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$, we obtain the desired result. ∎
Appendix E Proof of Theorem 5.5
We first restate Theorem 5.5 as follows:
Theorem E.1.
Suppose Lemma 5.4 hold, then we have
$\hat{\tau}(\mathbf{x})-\tau(\mathbf{x})=O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$ (32)
Proof.
As stated in Lemma 5.4, we have
$\hat{f}(\mathbf{x})-f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right),$ (33)
and thus
$\hat{f}^{\mu}(\mathbf{x})-f^{\mu}(\mathbf{x})=O_{p}\left((\frac{\mu}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{\mu}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{\mu}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$ (34)
Under Assumption 5.2, we have
$\hat{\mu}_{Y}^{O}(a,\mathbf{x})-\mu_{Y}^{O}(a,\mathbf{x})=O_{p}(n^{-\frac{\gamma}{2\gamma+d}})$ (35)
and
$\hat{\mu}_{S_{T}}^{E}(a,\mathbf{x})-\mu_{S_{T}}^{E}(a,\mathbf{x})=O_{p}(n^{-\frac{\alpha}{2\alpha+d}})$ (36)
and
$\hat{\mu}_{S_{T}}^{O}(a,\mathbf{x})-\mu_{S_{T}}^{O}(a,\mathbf{x})=O_{p}(n^{-\frac{\beta}{2\beta+d}}).$ (37)
Then, let $a:=f^{\mu}(\mathbf{X})$,$\hat{a}=\hat{f}^{\mu}(\mathbf{X})$, $b=\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})$, and similarly for $\hat{b}$, and under Assumption 5.3, $a$, $b$, $\hat{a}$ and $\hat{b}$ are all bounded. We analyze the term
$(ab-\hat{a}\hat{b})^{2}=$ $(ab-a\hat{b}+a\hat{b}-\hat{a}\hat{b})^{2}$ (38)
$=$ $(a(b-\hat{b})+(a-\hat{a})\hat{b})^{2}$
$\overset{(a)}{\leq}$ $2a^{2}(b-\hat{b})^{2}+2\hat{b}^{2}(a-\hat{a})^{2}$
$\overset{(a)}{\asymp}$ $(b-\hat{b})^{2}+(a-\hat{a})^{2},$
where the inequality $(a)$ is based on $(a+b)^{2}\leq 2(a^{2}+b^{2})$ and $(b)$ is based on the boundedness assumption. Then we have
$ab-\hat{a}\hat{b}=O_{p}\left((\frac{\mu}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{\mu}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{\mu}{(T-1)n})^{\frac{\beta}{2\beta+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}\right).$ (39)
Hence, the result is immediately proved by the form of $\tau(x)$ as $\tau(x)=\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{X})\left(\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})\right)$ and the result of Lemma 5.4. ∎

### v2

@@
 Abstract
-Long-term causal inference is an important but challenging problem across various scientific domains.
+Long-term causal inference is an important but challenging problem across various domains.
 To solve the latent confounding problem in long-term observational studies, existing methods leverage short-term experimental data.
 Ghassami et al.
-(2022) propose an approach based on the Conditional Additive Equi-Confounding Bias (CAECB) assumption, which asserts that the confounding bias in the short-term outcome is equal to that in the long-term outcome, so that the long-term confounding bias and the causal effects can be identified.
-While effective in certain cases, this assumption is limited to scenarios with a one-dimensional short-term outcome.
+[14] propose an approach based on the Conditional Additive Equi-Confounding Bias (CAECB) assumption, which asserts that the confounding bias in the short-term outcome is equal to that in the long-term outcome, so that the long-term confounding bias and the causal effects can be identified.
+While effective in certain cases, this assumption is limited to scenarios where there is only one short-term outcome with the same scale as the long-term outcome.
 In this paper, we introduce a novel assumption that extends the CAECB assumption to accommodate temporal short-term outcomes.
@@
 1 Introduction
-Long-term causal inference is an important but challenging problem across various scientific fields, such as education (Athey et al., 2019), medicine (Fleming et al., 1994), and marketing (Hohnhold et al., 2015).
+Long-term causal inference is an important but challenging problem across various scientific fields, such as education [4], medicine [11], and marketing [17].
 While in many real-world scenarios, long-term observational data are readily available, a major challenge for long-term causal inference is the presence of latent confounding in observational studies.
-A common way to mitigate this issue is incorporating short-term experimental data, which raises a fundamental question: how can short-term experimental data be leveraged to address latent confounding in observational data for long-term causal inference, as shown in Figure 1?
-Figure 1: Causal graphs of experimental data and observational data with $\mathbf{X}$ being covariates, $\mathbf{U}$ being latent confounders, $A$ being treatment, $\mathbf{S}$ being short-term outcome, and $Y$ being long-term outcome.
-Fig.
-1(a) represents the causal graph of the short-term experimental data, where treatment $A$ is not affected by latent confounders $\mathbf{U}$ and the long-term outcome $Y$ is unobserved.
-Fig.
-1(b) represents the causal graph of the long-term observational data, where the latent confounders $\mathbf{U}$ affect treatments $A$ and outcomes $\mathbf{S},Y$ and the long-term outcome $Y$ can be observed.
-Fig.
-1(c) represents the overall graph with expanding temporal short-term outcomes.
-Existing works explore various methods to mitigate latent confounding by utilizing observational and experimental data based on different assumptions.
-One widely used assumption is the Latent Unconfoundedness assumption (Athey et al., 2020; Chen & Ritzwoller, 2023; Yang et al., 2024), which posits the short-term potential outcomes mediate the long-term potential outcome in the observational data, i.e., $Y(a){\perp\!\!\!\perp}A|\mathbf{S}(a),\mathbf{X},G=O$.
-These methods are effective when short-term outcomes $\mathbf{S}$ contain substantial information about the latent confounder $\mathbf{U}$, such as in studies on the lifetime effects of youth employment and training programs in the United States (Aizer et al., 2024).
+A common way to mitigate this issue is to incorporate short-term experimental data, which raises a fundamental question: how can short-term experimental data be leveraged to address latent confounding in observational data for long-term causal inference, as shown in Fig.
+1(a) and 1(b)?
+Existing works explore various methods to mitigate latent confounding by utilizing observational and experimental data based on different assumptions  [^1] Please kindly refer to Appendix A for detailed discussion on more existing assumptions..
+One widely used assumption is the Latent Unconfoundedness assumption [5, 8, 28], which posits the short-term potential outcomes mediate the long-term potential outcome in the observational data, i.e., $Y(a){\perp\!\!\!\perp}A\mid\mathbf{S}(a),\mathbf{X},G=O$.
+These methods are effective when short-term outcomes $\mathbf{S}$ contain substantial information about the latent confounder $\mathbf{U}$, such as in studies on the lifetime effects of youth employment and training programs in the United States [1].
 However, this assumption essentially restricts that the latent confounders can not affect the long-term outcome, i.e., ruling out the causal edge $\mathbf{U}\rightarrow Y$.
-To address this limitation, a follow-up work (Ghassami et al., 2022) proposes a novel assumption called Conditional Additive Equi-Confounding Bias (CAECB), allowing for the causal link $\mathbf{U}\rightarrow Y$.
+To address this limitation, a follow-up work [14] proposes a novel assumption called Conditional Additive Equi-Confounding Bias (CAECB), allowing for the causal link $\mathbf{U}\rightarrow Y$.
 The CAECB assumption states that the confounding bias in the short-term outcome is equal to that in the long-term outcome, enabling the identification of long-term causal effects.
-This assumption may be more reasonable and more aligned with practical settings where confounding biases share a similar manner over short- and long-term outcomes conditional on covariates $\mathbf{X}$, such as in research studying the impact of school finance reforms on student outcomes (Jackson, 2016).
+This assumption may be more reasonable and more aligned with practical settings where confounding biases share a similar manner over short- and long-term outcomes conditional on covariates $\mathbf{X}$, such as in research studying the impact of school finance reforms on student outcomes [21].
 Following the work of Ghassami et al.
-(2022), we focus on the setting where the causal link $\mathbf{U}\rightarrow Y$ exists.
-Beyond this, we further consider temporal short-term outcomes, as shown in Figure 1(c), a more common scenario where the method under the CAECB assumption (Ghassami et al., 2022) is not applicable.
+[14], we focus on the setting where the causal link $\mathbf{U}\rightarrow Y$ exists.
+Beyond this, we further consider temporal short-term outcomes, as shown in Figure 1(c), a more common scenario where the method under the CAECB assumption [14] is not applicable.
 In many real-world applications, short-term outcomes exhibit temporal dependencies, and capturing these relationships is essential for inferring long-term causal effects.
@@
 Consequently, the absence of temporal considerations in existing methods limits their effectiveness and constrains the potential of long-term causal inference in practice.
+Figure 1: Causal graphs for experimental and observational data with $\mathbf{X}$ as covariates, $\mathbf{U}$ as latent confounders, $A$ as treatment, $\mathbf{S}$ as short-term outcomes, and $Y$ as the long-term outcome.
+Fig.
+1(a) shows the experimental data, where $A$ is unaffected by $\mathbf{U}$ and $Y$ cannot be unobserved.
+Fig.
+1(b) depicts the observational data, where $\mathbf{U}$ affects $A$, $\mathbf{S}$, and $Y$, and $Y$ can be observed.
+Fig.
+1(c) illustrates the full causal graph with temporally extended short-term outcomes.
 In this paper, we introduce a novel assumption called Functional CAECB (FCAECB), which extends the existing CAECB assumption to capture the temporal dependencies among temporal short-term outcomes.
@@
 2 Related Work
-Long-term causal inference For years, researchers have investigated which short-term outcomes can reliably predict long-term causal effects.
-Various criteria have been proposed for identifying valid surrogates, including the Prentice criteria (Prentice, 1989), principal surrogacy (Frangakis & Rubin, 2002), strong surrogate criteria (Lauritzen et al., 2004), causal effect predictiveness (Gilbert & Hudgens, 2008), and consistent surrogate and its variations (Chen et al., 2007; Ju & Geng, 2010; Yin et al., 2020).
+Long-term causal inference For years, researchers have investigated which short-term outcomes can reliably predict long-term causal effects [30, 12, 27, 15, 7, 22, 42].
 Recently, there has been growing interest in estimating long-term causal effects using surrogates, which is also the focus of this paper.
 One line of work assumes the unconfoundedness assumption.
-Under the unconfoundedness assumption, LTEE (Cheng et al., 2021) and Laser (Cai et al., 2024) are based on specifically designed neural networks for long-term causal inference.
-EETE (Kallus & Mao, 2020) explores the data efficiency from the surrogate in several settings and proposes an efficient estimator for treatment effect.
-ORL (Tran et al., 2023) introduces a doubly robust estimator for average treatment effects using only short-term experiments, additionally assuming stationarity conditions between short-term and long-term outcomes.
-Singh & Sun (2024) proposes a kernel ridge regression-based estimator for long-term effect under continuous treatment.
-Additionally, Wu et al.
-(2024) develop a policy learning method for balancing short-term and long-term rewards.
-Our work is different from them.
-We do not assume the unconfoundedness assumption, and we use the data combination technique to solve the problem of unobserved confounders.
-Another line of research, which avoids the unconfoundedness assumption, tackles the latent confounding problem by combining experimental and observational data.
-This setting is first introduced by the method proposed by Athey et al.
-(2019), which, under surrogacy assumption, constructs the so-called Surrogate Index (SInd) as the substitutions for long-term outcomes in the experimental data for effect identification.
+Under the unconfoundedness assumption, LTEE [9] and Laser [6] are based on specifically designed neural networks for long-term causal inference.
+EETE [23] explores the data efficiency from the surrogate in several settings and proposes an efficient estimator for treatment effects.
+ORL [37] introduces a doubly robust estimator for average treatment effects using only short-term experiments, additionally assuming stationarity conditions between short-term and long-term outcomes.
+Singh and Sun [34] proposes a kernel ridge regression-based estimator for long-term effect under continuous treatment.
+Additionally, Yang et al.
+[40], Wu et al.
+[39] develop a policy learning method for balancing short-term and long-term rewards.
+Our work is different from theirs.
+We do not rely on the unconfoundedness assumption, and we use the data combination technique to solve the problem of unobserved confounders.
+Another line of work, which avoids the unconfoundedness assumption, tackles the latent confounding problem by combining experimental and observational data.
+This setting is first introduced by Athey et al.
+[4], which, under surrogacy assumption, constructs the so-called Surrogate Index (SInd) as the substitutions for long-term outcomes in the experimental data for effect identification.
 As follow-up work, Athey et al.
-(2020) introduces the latent unconfoundedness assumption, which assumes that short-term potential outcomes can mediate the long-term potential outcomes, thereby enabling long-term causal effect identification.
-Under this assumption, Yang et al.
-(2024); Chen & Ritzwoller (2023) propose several estimators for effect estimation.
-The alternative feasible assumptions (Ghassami et al., 2022) are proposed to replace the latent unconfoundedness assumption, e.g., the Conditional Additive Equi-Confounding Bias (CAECB) assumption.
-Based on proximal methods (Tchetgen Tchetgen et al., 2024), Imbens et al.
-(2024) proposes considering the short-term outcomes as proxies of latent confounders, thereby achieving effect identification.
-However, these studies primarily concentrate on average treatment effects, whereas our focus in this paper is on heterogeneous effects.
-Among the existing literature, the most closely related work is Ghassami et al.
-(2022)’s work.
-Our work can be viewed as a significant extension of theirs, as we consider a more practical scenario where short-term outcomes exhibit temporal dependencies, and theoretically, their CAECB assumption is a special case of the assumption we propose.
+[5] introduce the latent unconfoundedness assumption, which assumes that short-term potential outcomes can mediate the long-term potential outcomes, thereby enabling long-term causal effect identification.
+Under this assumption, Meza and Singh [28], Yang et al.
+[41], Chen and Ritzwoller [8] propose several estimators for effect estimation.
+The alternative feasible assumptions [14] are proposed to replace the latent unconfoundedness assumption, e.g., the Conditional Additive Equi-Confounding Bias (CAECB) assumption.
+Based on proximal methods [36], Imbens et al.
+[19] propose considering the short-term outcomes as proxies of latent confounders, thereby achieving effect identification.
+Among the existing literature, the most closely related work is the work of Ghassami et al.
+[14].
+Our work can be viewed as a significant extension of theirs, as we consider a more practical scenario where short-term outcomes exhibit temporal dependencies, and theoretically, their CAECB assumption is a special case of our proposed assumption.
+Our work is also closely related to the work of Imbens et al.
+[19], which also explores the sequential short-term outcomes and assumes them as proxies of latent confounders.
+Different from that, we explore the sequential information from short-term confounding bias, which allows for inferring long-term confounding bias and results in the final effect identification.
 Modeling Latent Confounding Bias An effective way to solve the latent confounding problem in the data combination setting is to model latent confounding bias.
 Kallus et al.
-(2018) proposes modeling confounding bias under a linearity assumption.
+[24] proposes modeling confounding bias under a linearity assumption.
 Hatt et al.
-(2022) introduce to model the nonlinear confounding bias using the representation learning technique.
-Wu & Yang (2022) propose the integrative R-learner via a regularization for the conditional effects and confounding bias with the Neyman orthogonality.
+[16] introduce to model the nonlinear confounding bias using the representation learning technique.
+Wu and Yang [38] propose the integrative R-learner via a regularization for the conditional effects and confounding bias with the Neyman orthogonality.
 Zhou et al.
-(2025) propose a two-stage representation learning strategy to model such a confounding bias.
+[43] propose a two-stage representation learning strategy to model such a confounding bias.
 Different from these works, we focus on the long-term causal inference setting, and rather than focus on how to model the confounding bias, we concentrate more on the relationship between sequential confounding biases.
 3 Preliminaries
-3.1 Notations, Problem Definition, Assumptions
+3.1 Notations, Problem Definition, Basic Assumptions
 Let $A\in\{0,1\}$ be the treatment variable, $\mathbf{X}\in{\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ be the observed covariates where $d$ is the dimension of $\mathbf{X}$, $\mathbf{U}\in{\mathcal{U}}\subseteq{\mathbb{R}}^{d_{U}}$ be the latent confounders where $d_{U}$ is the dimension of $\mathbf{U}$.
 Let $\mathbf{S}=[S_{1},S_{2},\dots,S_{T}]$ be the short-term outcome variable where $S_{t}\in{\mathcal{S}}\subseteq{\mathbb{R}}$ is the short-term outcome measured at time step $t$, and $Y=S_{T+\mu}\in{\mathcal{R}}\subseteq{\mathbb{R}}$ be the long-term outcome.
-Further, we leverage the potential outcome framework proposed by Rubin (1978).
+Further, we leverage the potential outcome framework [32].
 We denote $\mathbf{S}(a)$ as the potential short-term outcome, $S_{t}(a)$ as the potential short-term outcome at time step $t$, and $Y(a)$ as the potential long-term outcome.
-Following existing work on long-term inference (Athey et al., 2020; Ghassami et al., 2022; Imbens et al., 2024; Chen & Ritzwoller, 2023), we denote $G\in\{E,O\}$ as the indicator of data group, where $G=E$ indicates experimental data and $G=O$ indicates observational data.
+Following existing work on long-term inference [5, 14, 19, 8], we denote $G\in\{E,O\}$ as the indicator of data group, where $G=E$ indicates experimental data and $G=O$ indicates observational data.
 Let lowercase letters (e.g., $a,\mathbf{x},\mathbf{u},\mathbf{s},y,s(a),y(a)$) denote the value of the previously described random variables.
@@
 For ease of convenience, we denote the following nuisance functions and confounding bias:
-$\mu_{S_{t}}^{E}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}|A,\mathbf{X},G=E],$ (1)
-$\mu_{S_{t}}^{O}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}|A,\mathbf{X},G=O],$
-$\mu_{Y}^{E}(A,\mathbf{X})={\mathbb{E}}[Y|A,\mathbf{X},G=E],$
-$\mu_{Y}^{O}(A,\mathbf{X})={\mathbb{E}}[Y|A,\mathbf{X},G=O],$
+$\mu_{S_{t}}^{E}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}\mid A,\mathbf{X},G=E],\quad\mu_{S_{t}}^{O}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}\mid A,\mathbf{X},G=O],$ (1)
+$\mu_{Y}^{E}(A,\mathbf{X})={\mathbb{E}}[Y\mid A,\mathbf{X},G=E],\quad\mu_{Y}^{O}(A,\mathbf{X})={\mathbb{E}}[Y\mid A,\mathbf{X},G=O],$
 $\omega_{t}(\mathbf{X})=\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X}),$
-where $\omega_{t}(\mathbf{X})$ is known as the confounding bias, the discrepancy between the conditional mean outcome differences derived from the experimental data and the observational data (see Sections 3.2 and 4 for more details on how it serves to the identification of long-term effects).
+where $\omega_{t}(\mathbf{X})$ is known as the confounding bias, the discrepancy between the conditional mean outcome differences derived from the experimental data and the observational data (Please see Sections 3.2 and 4 for more details on how it serves to the identification of long-term effects).
 Moreover, we denote stochastic boundedness with $O_{p}$ and convergence in probability with $o_{p}$.
@@
 We use $a_{n}\lesssim b_{n}$ to denote both $a_{n}\leq Cb_{n}$ for some constant $C>0$.
-Task: Given a short-term experimental dataset $\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},g^{(i)}=E\}_{i=1}^{n_{e}}$ and a long-term observational dataset $\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},y^{(i)},g^{(i)}=O\}_{i=n_{e}+1}^{n_{e}+n_{o}}$, the goal is to identify the following causal estimand of interest: the Heterogeneous Long-term Causal Effects (HLCE), i.e.,
-$\tau(\mathbf{x})={\mathbb{E}}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}].$ (2)
+Task: Given a short-term experimental dataset $\mathbb{D}_{e}$ and a long-term observational dataset $\mathbb{D}_{o}$, the goal in this paper is to identify and estimate the Heterogeneous Long-term Causal Effects (HLCE), i.e.,
+$\tau(\mathbf{x})={\mathbb{E}}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}].$ (2)
 However, HLCE $\tau(\mathbf{x})$ is not identifiable from the experimental data alone, since $Y$ is missing in that dataset.
-Also it is not identifiable from the observational data alone, since the observational data suffers from the latent confounding problem.
+Also it is not identifiable from the observational data alone, since the observational data suffers from the latent confounding problem, i.e., $A\not{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}\mid\mathbf{X},G=O$.
 Furthermore, the information regarding the causal effects in the experimental data is not necessarily relevant to that in the observational data without further assumptions.
-To ensure the identification, we first make the following assumptions that are commonly assumed in long-term causal inference:
+To ensure the identification, we first make the following assumptions that are commonly assumed in long-term causal inference [4, 5, 8, 14, 19]:
 Assumption 3.1 (Consistency).
-If a unit is assigned treatment, we observe its associated potential outcome. Formally, if $A=a$, then $Y=Y(a)$, and $\mathbf{S}=\mathbf{S}(a)$.
+If $A=a$, then $Y=Y(a)$, and $\mathbf{S}=\mathbf{S}(a)$.
 Assumption 3.2 (Positivity).
-The treatment assignment is non-deterministic. Formally, $\forall a,\mathbf{x}$, we have $10<P(A=a|\mathbf{X}=\mathbf{x})<$ 1, and $0<P(G=O|A=a,\mathbf{X}=\mathbf{x})<1$.
-Assumption 3.3 (Weak internal validity of observational data).
-Latent confounders exist in observational data.
-Formally, afor-$\forall a$, we have $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},\mathbf{U},G=O$ and $A\not{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},G=O$.
-Assumption 3.4 (Internal validity of experimental data).
-There are no latent confounders in experimental data. Formally, afor-$\forall a$, we have $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},G=E$.
-Assumption 3.5 (External validity of experimental data).
-The distribution of the potential outcomes is invariant to whether the data belongs to the experimental or observational data.
-Formally, afor-$\forall a$, we have $G{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X}$.
-Assumptions 3.1 and 3.2 are standard assumptions in causal inference (Rubin, 1974; Imbens, 2000).
-Assumptions 3.3, 3.4 and 3.5 are mild and widely assumed in data combination settings (Shi et al., 2023; Imbens et al., 2024; Athey et al., 2019, 2020; Hu et al., ).
+$\forall a,\mathbf{x}$, $0<P(A=a\mid\mathbf{X}=\mathbf{x});P(G=O\mid A=a,\mathbf{X}=\mathbf{x})<1$.
+Assumption 3.3 (Observational Data).
+afor-$\forall a$, $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}\mid\mathbf{X},\mathbf{U},G=O$.
+Assumption 3.4 (Experimental Data).
+afor-$\forall a$, $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}\mid\mathbf{X},G=E$.
+Assumption 3.5 (Data Combination).
+afor-$\forall a$, $G{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}\mid\mathbf{X}$.
+Assumptions 3.1 and 3.2 are standard assumptions in causal inference [31, 20].
+Assumptions 3.3, 3.4 and 3.5 are mild and widely assumed in data combination settings [33, 19, 4, 5, 18].
 Specifically, Assumption 3.3 allows the existence of latent confounders in observational data, thus it is much weaker than the traditional unconfoundedness assumption.
@@
 Assumption 3.5 connects the potential outcome distributions between observational and experimental data.
-Assumptions 3.1, 3.2, 3.3, 3.4 and 3.5 are still not sufficient to identify the causal estimand of interest.
+However, Assumptions 3.1, 3.2, 3.3, 3.4 and 3.5 are still not sufficient to identify the causal estimand of interest.
 The root cause is that, even though the assumptions above link the experimental and observational data, the (long-term) latent confounding problem remains unsolved.
 In the following section, we first review a method proposed by Ghassami et al.
-(2022), which poses an extra assumption, called Conditional Additive Equi-Confounding Bias (CAECB) assumption, to achieve the identification of HLCE.
+[14], which poses an extra assumption, called the Conditional Additive Equi-Confounding Bias (CAECB) assumption, to achieve the identification of HLCE.
 Then, we propose our approach that generalizes the method of Ghassami et al.
-(2022) to allow temporal short-term outcomes.
-(a) Schematic representation of CAECB assumption (Ghassami et al., 2022).
+[14] to allow temporal short-term outcomes.
+(a) Schematic representation of CAECB assumption [14].
 (b) Schematic representation of our proposed FCAECB assumption.
 Figure 2: Schematic representations of CAECB assumption proposed by Ghassami et al.
-(2022) and our FCAECB assumption.
+[14] and our FCAECB assumption.
 As shown in Figure 2(a), the CAECB assumption requires that the confounding bias in the short-term outcome is equal to that in the long-term outcome.
 As shown in Figure 2(b), the FCAECB assumption relaxes this constraint by allowing for temporal short-term outcomes and only requiring that confounding biases across different time steps follow a specific pattern rather than remaining equal.
-3.2 CAECB Assumption (Ghassami et al., 2022)
+3.2 CAECB Assumption [14]
 Ghassami et al.
-(2022) introduced a method for combining experimental and observational data to identify the HLCE under an extra CAECB assumption.
+[14] introduced a method for combining experimental and observational data to identify the HLCE under an extra CAECB assumption.
 The method under CAECB assumption is only applicable to scenarios with a one-dimensional short-term outcome, which we denote as $S$ (and its corresponding potential outcome $S(a)$) with slightly abusing notation.
-To begin with, we first restate the CAECB assumption as follows:
-Assumption 3.6 (Conditional Additive Equi-Confounding Bias, CAECB (Ghassami et al., 2022)).
+To begin with, we first restate the CAECB assumption:
+Assumption 3.6 (Conditional Additive Equi-Confounding Bias, CAECB [14]).
 The difference in conditional expected values of short-term potential outcomes across treated and control groups is the same as that of the long-term potential outcome variable.
 Formally, afor-$\forall a$, we have
-[…]
+${\mathbb{E}}[S(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[S(a)\mid\mathbf{X},A=1,G=O]$ (3)
+$=$ ${\mathbb{E}}[Y(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[Y(a)\mid\mathbf{X},A=1,G=O].$
 Theorem 3.7.
-Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6 hold, then $\tau(\mathbf{x})$ can be identified as follows:
-$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$ (4)
-$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+\omega(\mathbf{x}),$
+Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6 hold, then $\tau(\mathbf{x})$ can be identified:
+$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]=\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+\omega(\mathbf{x}),$ (4)
 where $\omega(\mathbf{x})=\mu_{S}^{E}(1,\mathbf{x})-\mu_{S}^{E}(0,\mathbf{x})+\mu_{S}^{O}(0,\mathbf{x})-\mu_{S}^{O}(1,\mathbf{x})$ is the short-term confounding bias.
-Proof can be found in Appendix A.
-A similar identification result in terms of long-term average causal effects has been shown by Ghassami et al.
-(2022).
-Here we provide the identification result of HLCE $\tau(\mathbf{x})$ in Theorem 3.7.
-Remark 3.8.
-We illustrate Assumption 3.6 in Figure 2(a).
-Assumption 3.6 can be seen as a generalization of the parallel assumption in the difference-in-differences (DiD) framework (Ashenfelter & Card, 1984; Angrist & Pischke, 2009), Note that Assumption 3.6 essentially implies that the short-term confounding bias is the same as the long-term one.
+Proof can be found in Appendix E.
+A similar identification result in terms of long-term average causal effects has been established by Ghassami et al.
+[14].
+For completeness, we present the identification result of the HLCE $\tau(\mathbf{x})$ in Theorem 3.7.
+Remark 3.8 (Interpretation on Assumption 3.6).
+Assumption 3.6 is intuitively illustrated in Figure 2(a).
+Assumption 3.6 can be seen as a generalization of the parallel assumption in the difference-in-differences (DiD) framework [2, 3], Note that Assumption 3.6 essentially implies that the short-term confounding bias is the same as the long-term one.
 In this way, $\omega(\mathbf{x})$ can be interpreted as the long-term confounding bias, resulting in the identification result in Eq.
@@
 Assumption 4.1 (Functional Conditional Additive Equi-Confounding Bias, FCAECB).
-The difference in conditional expected values of short-term potential outcomes across treated and control groups between two timesteps follows the learnable function forms $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$.
+The difference in conditional expected values of short-term potential outcomes across treated and control groups between two time steps follows the learnable function form $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$.
 Formally, afor-$\forall a$, we have
 $b_{t+1}(a,\mathbf{X})=f(\mathbf{X})b_{t}(a,\mathbf{X}),$ (5)
-where $b_{t}(a,\mathbf{X})={\mathbb{E}}[S_{t}(a)|\mathbf{X},A=0,G=O]-{\mathbb{E}}[S_{t}(a)|\mathbf{X},A=1,G=O]$.
-Remark 4.2.
+where $b_{t}(a,\mathbf{X})={\mathbb{E}}[S_{t}(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[S_{t}(a)\mid\mathbf{X},A=1,G=O]$.
+Remark 4.2 (Interpretation on Assumption 4.1).
 We illustrate our proposed FCAECB assumption in Figure 2(b).
-We relax the CAECB assumption (Ghassami et al., 2022) to allow for the temporal short-term outcomes, instead of restricting that the short-term outcome should be the same scale as the long-term outcome.
-Additionally, the existing CAECB assumption can be seen as our special case when $f(\mathbf{X})$ in the FCEACB assumption satisfies $f(\mathbf{X})=1$.
+We relax the CAECB assumption [14] to allow for the temporal short-term outcomes, instead of restricting that the short-term outcome should be the same scale as the long-term outcome.
+Additionally, the existing CAECB assumption can be seen as our special case when $f(\mathbf{X})$ in the FCAECB assumption satisfies $f(\mathbf{X})=1$.
 To provide a better understanding of our Assumption 4.1, we provide the insight in term of the functional form $f(\mathbf{X})$ in the following proposition.
@@
 where $\omega_{t}(\mathbf{X})$ is the confounding bias at time step $t$, defined as $\omega_{t}(\mathbf{X})=\mu_{S_{t}}^{E}(1,\mathbf{X})-\mu_{S_{t}}^{E}(0,\mathbf{X})+\mu_{S_{t}}^{O}(0,\mathbf{X})-\mu_{S_{t}}^{O}(1,\mathbf{X})$.
-The proof is given in Appendix B.
-Remark 4.4.
+The proof is given in Appendix F.
+Remark 4.4 (Insight of Assumption 4.7).
 As stated in Proposition 4.3, Assumption 4.1 essentially states the confounding biases $\omega_{t}(\mathbf{X})$ between adjacent time steps follow the functional form $f(\mathbf{X})$.
@@
 (5) defining $f(\mathbf{X})$ using the potential outcomes (See Section 5 for the estimation based on Proposition 4.3).
+Remark 4.5 (Testing Assumption 4.7 over Short Term Duration).
+It is possible to empirically test Assumption 4.1 over the short-term duration, which provides insights into whether it can be extrapolated to long-term confounding bias.
+This is particularly reasonable when short- and long-term outcomes share the same practical meaning, making it more likely that the confounding bias mechanism remains stable.
+Specifically, suppose the short-term outcomes are observed at $3$ time steps.
+Then, given $S_{1},S_{2},S_{3}$, we can estimate $f(\mathbf{x})$ using $S_{1}$ and $S_{2}$, and evaluate the $R^{2}$ measure on $S_{2}$ and $S_{3}$ where $R^{2}=1-\frac{\sum_{i}^{N}(\omega_{3}(\mathbf{x}^{(i)})-\hat{f}(\mathbf{x}^{(i)})\omega_{2}(\mathbf{x}^{(i)}))^{2}}{\sum_{i}^{N}(\omega_{3}(\mathbf{x}^{(i)})-\frac{1}{N}\sum_{j}^{N}\omega_{3}(\mathbf{x}^{(j)}))^{2}}$.
+If $R^{2}\approx 1$, it suggests that $f$ well explains the relationships of confounding biases between adjacent time steps.
+Conversely, if $R^{2}\ll 1$, then the assumption may not hold, or $f$ is poorly fitted.
+Example 4.6.
+In many real-world scenarios, Assumption 4.1 is reasonable.
+In the motivating example from our introduction, AIDs patients undergo regular follow-up visits, and their short-term health indicators, such as viral loads or CD4 counts, are recorded as short-term outcomes.
+Since CD4 counts are critical indicators of health status, they serve as both short- and long-term outcomes.
+In such cases, Assumption 4.1 can be tested over short-term durations and is likely to hold.
 The key to identifying long-term causal effect under Assumption 4.1 is using the temporal information in $f(\mathbf{X})$ to extrapolate the long-term confounding bias $\omega_{T+\mu}(\mathbf{X})$.
@@
 Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 4.1 hold, then the heterogeneous long-term effects $\tau(\mathbf{x})$ can be identified as follows:
-$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$ (7)
-$=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x}),$
+$\tau(\mathbf{x})=$ $\mathbb{E}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]=\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{x})\omega_{T}(\mathbf{x}),$ (7)
 where $\omega_{T}(\mathbf{x})=\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})$ is the short-term confounding bias at time step $T$.
-The proof is given in Appendix C.
+The proof is given in Appendix G.
 Theorem 4.7 provides the identification result of the heterogeneous long-term effects $\tau(\mathbf{X})$.
@@
   • S3.
-    Constructing final HLCE estimator as
-    $\hat{\tau}(\mathbf{x})=\hat{\mu}_{Y}^{O}(1,\mathbf{x})-\hat{\mu}_{Y}^{O}(0,\mathbf{x})+\hat{f}^{\mu}(\mathbf{x})$ (9)
-    ×$\times\left(\hat{\mu}_{S_{T}}^{E}(1,\mathbf{x})-\hat{\mu}_{S_{T}}^{E}(0,\mathbf{x})+\hat{\mu}_{S_{T}}^{O}(0,\mathbf{x})-\hat{\mu}_{S_{T}}^{O}(1,\mathbf{x})\right).$
-Note that, in addition to the fitting/constructing steps S1-S3, we also introduce an optional S0 to select subsets of short-term outcomes $\mathbf{S}$.
+    Constructing the final HLCE estimator as
+    $\hat{\tau}(\mathbf{x})=\hat{\mu}_{Y}^{O}(1,\mathbf{x})-\hat{\mu}_{Y}^{O}(0,\mathbf{x})+\hat{f}^{\mu}(\mathbf{x})\left(\hat{\mu}_{S_{T}}^{E}(1,\mathbf{x})-\hat{\mu}_{S_{T}}^{E}(0,\mathbf{x})+\hat{\mu}_{S_{T}}^{O}(0,\mathbf{x})-\hat{\mu}_{S_{T}}^{O}(1,\mathbf{x})\right).$ (9)
+Note that, in addition to the fitting/constructing steps S1-S3, we also introduce an optional step S0 to select appropriate subsets of short-term outcomes $\mathbf{S}$.
 This step is motivated by the identification result in Eq.
@@
 5.2 Convergence Rate Analyses
-In this paper, we assume the smoothness assumption of the estimated functions, where $s$-smooth functions are contained in the Hölder ball $\mathcal{H}_{d}(s)$, estimable with the minimax rate (Stone, 1980) of $n^{\frac{1}{2+d/s}}$ where $d$ is the dimension of $\mathcal{X}$.
+In this paper, we assume the smoothness assumption of the estimated functions, where $s$-smooth functions are contained in the Hölder ball $\mathcal{H}_{d}(s)$, estimable with the minimax rate [35] of $n^{\frac{1}{2+d/s}}$ where $d$ is the dimension of $\mathcal{X}$.
 Formally, we provide the following definition.
@@
 $\forall x,x^{\prime}$ and $m=(m1,\cdots,m_{d})$ such that $\Sigma_{j=1}^{d}m_{j}=\lfloor s\rfloor$.
+Accordingly, we assume the smoothness assumption of the nuisance functions, as well as the boundedness assumption of the nuisance functions and their estimates as follows.
 Assumption 5.2 (Smoothness Assumption).
-We assume that the nuisance functions and the mapping $f$ defined in Assumption 4.1 satisfy: (1) $\mu_{S_{t}}^{E}$, $\mu_{S_{t}}^{O}$, and $\mu_{Y}^{O}$ are $\alpha$-smooth, $\beta$-smooth, and $\gamma$-smooth, respectively, and all are estimable at Stone (1980)’s minimax rate of $n^{\frac{-p}{2p+d}}$ for a $p$-smooth function; (2) $f(x)$ is $\eta$-smooth.
-We also assume the boundedness assumption of the nuisance functions.
+We assume that the nuisance functions and the mapping $f$ defined in Assumption 4.1 satisfy: (1) $\mu_{S_{t}}^{E}$, $\mu_{S_{t}}^{O}$, and $\mu_{Y}^{O}$ are $\alpha$-smooth, $\beta$-smooth, and $\gamma$-smooth, respectively, and all are estimable at Stone [35]’s minimax rate of $n^{\frac{-p}{2p+d}}$ for a $p$-smooth function; (2) $f(x)$ is $\eta$-smooth.
 Assumption 5.3 (Boundness Assumption).
@@
 We now state our main theoretical results.
-To obtain our error bounds of the proposed estimator, we leverage the same sample splitting technique from (Kennedy, 2023), which randomly splits the datasets into two independent sets and applies them to the regressions of the first step and second step respectively.
-Such a technique is originally used to analyze the convergence rate of the double robust conditional average treatment effect estimation in the traditional setting (Kennedy, 2023) and later is adapted to several other methods (Curth & Van der Schaar, 2021; Frauen & Feuerriegel, ).
+To obtain our error bounds of the proposed estimator, we leverage the same sample splitting technique from [25], which randomly splits the datasets into two independent sets and applies them to the regressions of the first step and second step respectively.
+Such a technique is originally used to analyze the convergence rate of the double robust conditional average treatment effect estimation in the traditional setting [25] and later is adapted to several other methods [10, 13].
 Different from them, we use such a technique for the sequential latent confounding modeling, which is then adapted for the long-term effect estimation.
@@
 Lemma 5.4.
-Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, and 5.2 hold, then we have
-$\hat{f}(\mathbf{x})-$ $f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}\right.$ (10)
-$\left.+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$
-The proof is given in Appendix D.
+Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, 5.2, and 5.3 hold, then we have
+$\hat{f}(\mathbf{x})-$ $f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right),$ (10)
+which attains the oracle rate if $\min\{\alpha,\beta\}\geq\eta$.
+The proof is given in Appendix H.
 Since $\hat{f}(\mathbf{x})$ is a time-series model, its effective sample is $(T-1)n$, thus it might achieve a faster rate if we observed a longer duration of short-term outcomes.
@@
 $\hat{\tau}(\mathbf{x})-\tau(\mathbf{x})$ (11)
-$=$ $O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}\right.$
-$\left.+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$
-Proof can be found in Appendix E.
+$=$ $O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$
+Proof can be found in Appendix I.
 Theorem 5.5 follows directly from Lemma 5.4.
-The rate of $\hat{\tau}(\mathbf{x})$ consist of several terms: 1.
-$n^{-\frac{\gamma}{2\gamma+d}}$ represents the rate of $\mu_{Y}^{O}(a,\mathbf{x})$, 2.
-$n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}$ corresponds the rate of confounding bias $\omega_{T}(\mathbf{x})$, 3.
-and the remaining term is the rate of $\hat{f}^{\mu}(\mathbf{x})$.
-Theorem 5.5 suggests that the convergence rate is primarily influenced by the smoothness parameters $\alpha,\beta,\gamma,\eta$ and the temporal parameters $T,\mu$, where $T$ represents the longest observed duration of the short-term outcomes $\mathbf{S}$ and $\mu$ represents the time horizon of the long-term outcome to be estimated.
-This implies that achieving a more accurate effect estimation requires observing short-term outcomes for as long as possible.
+The rate of $\hat{\tau}(\mathbf{x})$ consist of three components: (a) $n^{-\frac{\gamma}{2\gamma+d}}$ represents the rate of $\mu_{Y}^{O}(a,\mathbf{x})$; (b) $n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}$ corresponds the rate of confounding bias $\omega_{T}(\mathbf{x})$; (c) and the remaining term is the rate of $\hat{f}^{\mu}(\mathbf{x})$.
+Theorem 5.5 indicates that the overall convergence rate is largely governed by the smoothness parameters $\alpha,\beta,\gamma,\eta$ and the temporal information $T$ and $\mu$, where $T$ denotes the maximum observed duration of short-term outcomes $\mathbf{S}$, and $\mu$ is the time horizon of the long-term outcome of interest.
+This result aligns with intuition: both $T$ and $\mu$ affect the convergence behavior.
+More accurate effect estimation benefits from longer short-term outcome durations $T$, while requiring the long-term horizon $\mu$ not to be too distant.
 Remark 5.6 (Choosing $\mu$ and $T$).
@@
 6.1 Experimental Setup
-In this section, we describe the data generation process for the synthetic datasets, the implementation details, and the metrics used in the experiments.
-Synthetic Datasets The data generation process is partly following (Kallus et al., 2018) such that we can obtain a specific form of confounding bias.
-Specifically, we first generate the treatments as follows: $A|G=O\sim\text{Bernoulli}(0.6)$ and $A|G=E\sim\text{Bernoulli}(0.4)$.
-Then we generate the observed $\mathbf{X}$ and the unobserved $\mathbf{U}$ as follows:
-$(\mathbf{X},\mathbf{U})|A,G=E\sim\mathcal{N}([\frac{2A-1}{2},0],\begin{bmatrix}1&0\\ 0&1\end{bmatrix})$ (12)
-$(\mathbf{X},\mathbf{U})|A,G=O\sim\mathcal{N}([\frac{1-2A}{2},0],\begin{bmatrix}1&A-0.5\\ A-0.5&1\end{bmatrix}).$
-Finally, we generate the $T$-step short-term outcomes $\mathbf{S}$ and the long-term outcome $Y=S_{\mu}$ satisfying Assumption 4.1 as follows:
-$S_{t}(1)=1+1.1\mathbf{X}+\mathbf{U}+\sum_{k=0}^{t}S_{k}(1)+\epsilon_{S_{t}},$ (13)
-$S_{t}(0)=\mathbf{X}+\mathbf{U}+\sum_{k=0}k^{t}S_{k}(0)+\epsilon_{S_{t}},$
-where $\epsilon_{S_{t}}$ are Gaussian noises.
-This will result in non-equal confounding bias in different time steps $t$, i.e., $\omega_{t}(\mathbf{x})=t\times\mathbf{x}$.
-We perform several control experiments using this data generation process.
-The default values are: the observational data sample size $n_{o}=4000$, experimental data sample size $n_{e}=2000$, $T=6$, and $\mu=3$.
-In control experiments, we vary $\mu$ within $\{1,2,3,4,5\}$ and $T$ within $T\in\{4,5,6,7,8\}$.
-We also fix the ratio of $n_{O}:n_{E}=2:1$, and varying $n_{e}$ within $\{1000,2000,3000,4000,5000,10000\}$ (which corresponds to $n_{o}$ values of $\{2000,4000,6000,8000,10000,20000\}$).
-Implementation We fit all nuisance functions using correctly specified regression.
-We denote our estimator as $\hat{\tau}$.
-For comparison, we evaluate our method against a correctly specified T-Learner (Künzel et al., 2019), denoted as $\hat{\tau}_{exp}$, which directly regresses unobserved $Y$ in the experimental data.
-However, this estimator $\hat{\tau}_{exp}$ is infeasible in practice since $Y$ is missing in the experimental data.
-This baseline serves as an idealized benchmark, and comparing our method against it highlights the effectiveness of our estimator across different control experiments.
+Table 1: The results on synthetic data are presented as the mean and standard deviation.
+Lower values indicate better performance.
+The best-performing methods are bolded, with the exception of the idealized estimator $\hat{\tau}_{exp}$, which is excluded from the comparison.
+(See Appendix D for Full results.)
+Data $\mu=1$, $T=6$ $\mu=3$, $T=6$ $\mu=5$, $T=6$
+Metrics ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$
+$\hat{\tau}_{exp}$ (idealized) 3.0033 $\pm$ 1.8226 2.0722 $\pm$ 1.7681 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 48.0590 $\pm$ 29.1580 33.1620 $\pm$ 28.2845
+T-learner (using obs. $Y$) 71.8431 $\pm$ 3.1612 6.8211 $\pm$ 2.8357 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3424 1149.4769 $\pm$ 50.5746 109.1205 $\pm$ 45.3685
+LTEE 73.0421 $\pm$ 7.1531 10.0977 $\pm$ 6.1071 287.3264 $\pm$ 28.5202 32.4417 $\pm$ 19.5739 1168.7612 $\pm$ 122.7687 137.0583 $\pm$ 107.5567
+Latent Unconfoundedness - 2.6322 $\pm$ 2.0120 - 10.5293 $\pm$ 8.0471 - 42.1179 $\pm$ 32.1874
+Sequential Outcomes ($S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$) - 67.5881 $\pm$ 3.0107 - 270.3607 $\pm$ 12.0428 - 1081.4361 $\pm$ 48.1759
+Sequential Outcomes ($S_{1},S_{\lceil T/2\rceil},S_{T}$) - 67.5790 $\pm$ 3.0138 - 270.3237 $\pm$ 12.0553 - 1081.2884 $\pm$ 48.2252
+Sequential Outcomes ($S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$) - 67.5797 $\pm$ 3.0151 - 270.3266 $\pm$ 12.0607 - 1081.3001 $\pm$ 48.2472
+NNPIV_LU 70.7124 $\pm$ 2.5320 6.1302 $\pm$ 2.8888 282.8595 $\pm$ 10.1320 24.5293 $\pm$ 11.5558 1131.4331 $\pm$ 40.5209 98.1118 $\pm$ 46.2263
+NNPIV_SInd 70.7155 $\pm$ 2.5329 6.1306 $\pm$ 2.8884 282.8655 $\pm$ 10.1314 24.5240 $\pm$ 11.5558 1131.4602 $\pm$ 40.5252 98.0989 $\pm$ 46.2232
+CAECB (using first $S$) 70.7257 $\pm$ 3.1153 6.7160 $\pm$ 2.7984 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 1148.3595 $\pm$ 50.5286 109.0154 $\pm$ 45.3308
+CAECB (using middle $S$) 67.3754 $\pm$ 2.9827 6.3982 $\pm$ 2.6884 282.9014 $\pm$ 12.4626 26.8566 $\pm$ 11.1910 1145.0089 $\pm$ 50.3933 108.6976 $\pm$ 45.2162
+CAECB (using last $S$) 36.1202 $\pm$ 2.0211 3.5778 $\pm$ 1.8603 251.6244 $\pm$ 11.2383 23.8972 $\pm$ 10.2153 1113.7291 $\pm$ 49.1287 105.7381 $\pm$ 44.1729
+CAECB (using random $S$) 59.8772 $\pm$ 11.9929 5.7194 $\pm$ 2.6871 275.4020 $\pm$ 16.0775 26.1778 $\pm$ 10.9897 1137.5094 $\pm$ 50.1542 108.0188 $\pm$ 44.9689
+Ours $\hat{\tau}$ 3.1110 $\pm$ 1.8230 2.1470 $\pm$ 1.7826 13.9945 $\pm$ 7.1682 9.296 $\pm$ 7.0820 64.3076 $\pm$ 32.9868 40.5425 $\pm$ 29.7264
+In this section, we describe the data generation process for the synthetic datasets, the baselines, the implementation details, and the metrics used in the experiments.
+Synthetic Datasets We partially follow the data generation setup in [24] to induce a specific form of confounding bias and define nuisance functions.
+Details are provided in Appendix D.
+The confounding bias at time $t$ is designed as $\omega_{t}(\mathbf{x})=-2^{t}\mathbf{x}$, aligning with our assumptions.
+By default, we set the observational sample size $n_{o}=4000$, experimental sample size $n_{e}=2000$, and $T=6$, $\mu=3$.
+In variations, we change $\mu\in{1,2,3,4,5}$, $T\in{4,5,6,7,8}$, and fix the ratio $n_{o}:n_{e}=2:1$ while varying $n_{e}\in\{1000,2000,3000,4000,5000,10000\}$ (implying $n_{o}\in\{2000,4000,6000,8000,10000,2000\}$).
+Baselines and Implementation We compare our method against several baselines for long-term causal effect estimation, including LTEE [9], Latent Unconfoundedness [5], Sequential Outcomes [20], NNPIV [28], and CAECB [14].
+For the Sequential Outcomes method, which requires three groups of short-term outcomes ($S_{1},S_{2},S_{3}$ in their paper), we test three grouping strategies: (1) $S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$, (2) $S_{1},S_{\lceil T/2\rceil},S_{T}$, and (3) $S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$.
+NNPIV is applied under both the latent unconfoundedness assumption (NNPIV_LU) [16] and the surrogacy assumption (NNPIV_SInd) [4].
+For CAECB, which uses a single short-term outcome, we implement four variants using the first, middle, last, and a randomly selected outcome.
+We also include the T-learner [26] trained on observational $Y$ and an idealized version trained on experimental $Y$ (infeasible in practice, as $Y$ is unobserved in experimental data), serving as an idealized benchmark denoted as $\hat{\tau}_{exp}$.
+All nuisance functions are estimated using correctly specified regression.
+Our estimator is denoted as $\hat{\tau}$.
+Additional details are provided in Appendix D.
 Metrics As for heterogeneous effect estimation, we report Precision in the Estimation of Heterogeneous Effect (PEHE) $\varepsilon_{PEHE}=\sqrt{\frac{1}{n}\Sigma_{i=1}^{n}(\tau(\mathbf{x}_{i})-\hat{\tau}(\mathbf{x}_{i}))^{2}}$.
 As for average long-term causal effect estimation, we report the absolute error $\varepsilon_{ATE}=|\Sigma_{i=1}^{n}\tau(\mathbf{x}_{i})-\Sigma_{i=1}^{n}\hat{\tau}(\mathbf{x}_{i}))|$.
-For all metrics, we report the mean values and standard errors by $100$ times running.
+For all metrics, we report the mean values and standard deviation (in tables) or standard errors (in figures) by $50$ running.
 6.2 Experimental Results
-(a) Heterogeneous effect estimation error.
-(b) Average effect estimation error.
+Comparison with Baselines We compare our method with existing methods using different choices of $T$ and $\mu$, and results are shown in Fig.
+1 As expected, our method consistently outperforms all baselines, primarily because it is built upon more appropriate assumptions.
+In contrast, the baseline methods suffer from substantial biases resulting from incorrect assumptions.
+Furthermore, we observe that our estimator closely matches the idealized estimator in both magnitude and trend, highlighting the effectiveness of the proposed method.
+Optimal Choice of $T$ and $\mu$ We conduct experiments using different choices of $T$ and $\mu$. Specifically, to predict long-term effects, we consider the following subsets of $\mathbf{S}$:
+(a) Result of $\varepsilon_{PEHE}$.
+(b) Result of $\varepsilon_{ATE}$.
 Figure 3: Results of the experiments in terms of different choice of $\mu$ and $T$.
-Optimal Choice of $T$ and $\mu$: We conduct experiments using different choice of $T$ and $\mu$.
-Specifically, to predict long-term effects, we consider the following subsets of $\mathbf{S}$: 1.
+1.
 $\mathbf{S}=[S_{1},S_{3}]$ with $T=2,\mu=3$; 2.
 $\mathbf{S}=[S_{1},S_{3},S_{5}]$ with $T=3,\mu=2$; 3.
-$\mathbf{S}=[S_{1},S_{2},S_{3},S_{4},S_{5}]$ with $T=5,\mu=3$; and 4.
+$\mathbf{S}=[S_{1},S_{2},S_{3},S_{4},S_{5},S_{6}]$ with $T=6,\mu=3$; and 4.
 $\mathbf{S}=[S_{1},S_{5}]$ with $T=2,\mu=1$.
@@
 This may be attributed to the fact that our estimator $\hat{\tau}(\mathbf{x})$ leverages observational data, leading to higher data efficiency.
-(a) Heterogeneous effect estimation error with fixed $T$ and varying $\mu$.
-(b) Average effect estimation error with fixed $T$ and varying $\mu$.
-(c) Heterogeneous effect estimation error with fixed $\mu$ and varying $T$.
-(d) Average effect estimation error with fixed $\mu$ and varying $T$.
+(a) Fix $T$ and vary $\mu$.
+(b) Fix $T$ and vary $\mu$.
+(c) Fix $\mu$ and vary $T$.
+(d) Fix $\mu$ and vary $T$.
 Figure 4: Results of control experiments with fixed $T$ and varying $\mu$ (Figures 4(a) and 4(b)), and with fixed $\mu$ and varying $T$ (Figures 4(c) and 4(d)).
-Vary $T$ and $\mu$: We conduct control experiments with varying $T$ and varying $\mu$ separately.
+Vary $T$ and $\mu$ We conduct control experiments with varying $T$ and varying $\mu$ separately.
 First, Figures 4(a) and 4(b)) show the results for the experiments with fixed $T$ and varying $\mu$.
@@
 Moreover, we observe that, with smaller $\mu$, our estimator closely matches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, showing the effectiveness of our estimator.
-(a) Heterogeneous effect estimation error.
-(b) Average effect estimation error.
+(a) Results of $\varepsilon_{PEHE}$.
+(b) Results of $\varepsilon_{ATE}$.
 Figure 5: Results of control experiments with varying sample sizes $n_{e}$ and $n_{o}$.
-Sample sensitivity: We conduct control experiments by varying the sample sizes of both experimental and observational data.
+Sample size sensitivity We conduct control experiments by varying the sample sizes of both experimental and observational data.
 The results are presented in Figure 5.
@@
 Notably, when $n_{o}$ and $n_{e}$ become larger, our estimator $\hat{\tau}(\mathbf{x})$ closely approaches the idealized estimator $\hat{\tau}_{exp}(\mathbf{x})$, which again confirms the effectiveness of our estimator.
-Furthermore, when the sample size $n_{e}\geq 3000$ ($n_{e}\geq 6000$), the estimation errors become stable, indicating an appropriate sample size threshold.
+Furthermore, when the sample size $n_{e}\geq 3000$ ($n_{o}\geq 6000$), the estimation errors become stable, indicating an appropriate sample size threshold.
 7 Conclusion
-In this paper, we extend the existing CAECB assumption (Ghassami et al., 2022) for long-term causal inference.
-The original assumption is restricted to settings with a one-dimensional short-term outcome.
-To address this limitation, we introduce a more general assumption—the functional CAECB assumption—which accommodates temporal short-term outcomes.
-We theoretically establish the identification result of the heterogeneous long-term causal effect under our assumption and propose a corresponding estimator by modeling sequential latent confounding.
-Additionally, we provide a comprehensive theoretical analysis of the estimator.
-Experiments confirm both the validity of our theoretical results and the effectiveness of the proposed estimator.
-Impact Statement
-This paper presents work whose goal is to advance the field of Long-term causal inference in the presence of latent confounders.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
-Appendix A Proof of Theorem 3.7
+In this paper, we extend the existing CAECB assumption [14] for long-term causal inference.
+The original assumption is restricted to settings where short- and long-term outcomes share the same scale.
+To address this limitation, we introduce a more general assumption, the functional CAECB assumption, which accommodates temporal short-term outcomes.
+Under this new assumption, we theoretically establish identification of heterogeneous long-term causal effects and practically develop an estimator by modeling sequential latent confounding.
+We further provide a comprehensive theoretical analysis of the estimator.
+Experimental results support both the theoretical findings and the estimator’s effectiveness.
+Supplement to
+“Long-term Causal Inference via Modeling Sequential Latent Confounding”
+Appendix organization:
+[…]
+1
+Appendix A Assumption Discussion: Latent Unconfoundedness [5], Sequential Outcomes [19], CAECB [14], FCAECB
+In this section, we compare different key assumptions in [5, 19, 14] and our FCAECB assumption, and show the differences in terms of causal graphs.
+Assumptions Statements Causal graphs
+Latent Unconfoundedness [5] $Y(a){\perp\!\!\!\perp}A\mid\mathbf{X},\mathbf{S}(a),G=O$ [Uncaptioned image]
+Sequential Outcomes [19] $(Y(a),\mathbf{S}_{3}(a)){\perp\!\!\!\perp}\mathbf{S}_{1}(a)\mid\mathbf{S}_{2}(a),\mathbf{U},\mathbf{X},G=O$ [Uncaptioned image]
+CAECB [14] ${\mathbb{E}}[S(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[S(a)\mid\mathbf{X},A=1,G=O]$ $={\mathbb{E}}[Y(a)\mid\mathbf{X},A=0,G=O]-{\mathbb{E}}[Y(a)\mid\mathbf{X},A=1,G=O]$ [Uncaptioned image]
+Ours FCAECB $b_{t+1}(a,\mathbf{X})=f(\mathbf{X})b_{t}(a,\mathbf{X})$ where $b_{t}(a,\mathbf{X})={\mathbb{E}}[S_{t}(a)\mid\mathbf{X},A=0,G=O]$                                $-{\mathbb{E}}[S_{t}(a)\mid\mathbf{X},A=1,G=O]$ [Uncaptioned image]
+Table 2: Various assumptions and corresponding causal graphs.
+The Latent Unconfoundedness assumption states that unobserved confounders affect the long-term outcome only through short-term outcomes.
+In terms of causal graphs, $\mathbf{S}(a)$ fully mediates the path from $A$ to $Y(a)$, which results in ruling out the causal edge from $\mathbf{U}$ to $Y$ (Also see the discussion in [14]).
+This assumption ensures that $\mathbf{S}$ can serve a role of $\mathbf{U}$, making the unconfoundedness assumption hold in a certain way and thus identifying long-term effects.
+Instead, the Sequential Outcomes assumption and CAECB assumption allow for such a causal edge $\mathbf{U}\rightarrow Y$, while they posit different requirements.
+The Sequential Outcomes assumption requires that the short- and long-term outcomes can only be directly affected by the outcomes immediately preceding them, but not outcomes further in the past.
+In other words, $\mathbf{S}_{t+1}$ can only be affected by $\mathbf{S}_{t}$, but not by $\mathbf{S}_{t-1}$ or earlier.
+This assumption ensures that $\mathbf{S}_{1}$ and $\mathbf{S}_{3}$ act similarly as negative controls in proximal causal learning [29, 36], making effects identifiable.
+The CAECB assumption does not posit any requirement on the causal graph, but requires that the short- and long-term outcomes are measured on the same scale.
+Thus, the short-term outcome becomes a scale variable and may be easily violated when many short-term outcomes can be observed.
+This assumption states that the confounding biases in short and long term are the same, making long-term confounding bias can be identified using the observed short-term outcomes.
+Our FCAECB assumption likewise places no graph restriction, and further, we allow for temporal short-term outcomes instead of a single $S$.
+The FCAECB assumption posits that the sequence of short-term confounding biases follows a learnable temporal trend, which enables us to infer the long-term bias from observed $\mathbf{S}_{1:T}$.
+Compared with the Latent Unconfoundedness and Sequential Outcomes assumptions, our proposed FCAECB assumption is neither logically stronger nor weaker.
+While those two assumptions impose graphical constraints, FCAECB instead posits a particular form of data-generating process.
+And compared to the CAECB assumption, the FCAECB assumption is weaker, as CAECB corresponds to a special case of FCAECB where $f(\mathbf{X})=1$.
+Our FCAECB assumption can be further relaxed by allowing it to vary over time.
+In other words, we can replace $f(\mathbf{X})$ with $f(\mathbf{X},t)$ without any other modification, and the identification result also holds with almost the same proofs.
+However, this generalization introduces a new challenge: how to extrapolate $f(\mathbf{X},t)$, learned with $t\in\{1,...,T\}$, to $f(\mathbf{X},T+\mu)$.
+Addressing this challenge requires additional assumptions on $f$, such as temporal smoothness, autoregressive dependencies, etc.
+We leave this extension for future work.
+Appendix B Broader impacts
+Our work focuses on long-term causal inference, whose goal is to estimate long-term causal effects by combining long-term observational data and short-term experimental data.
+Our proposed method could be applied to a wide range of applications, such as decision-making in marketing and preventive measure studies in epidemiology.
+There are no specific possible harms or negative societal impacts that we should claim.
+Appendix C Limitations
+While our work relaxes some of the existing assumptions, it still relies on relatively strong ones.
+Regarding data requirements, similar to [5, 14, 19], our work depends on access to short-term experimental data, which may not be feasible in certain contexts, such as studies on smoking where interventions would be unethical.
+Additionally, the core Assumption 4.1 proposed in this paper may not hold in practice and remains untestable due to the absence of long-term outcomes in experimental data, although we propose a partial validation using short-term outcomes.
+Appendix D Additional Experimental Details and Results
+Data Generation Process Our data generation process partly follows [24].
+Specifically, we first generate the treatments as follows: $A\mid G=O\sim\text{Bernoulli}(0.6)$ and $A\mid G=E\sim\text{Bernoulli}(0.4)$.
+Then we generate the observed $\mathbf{X}$ and the unobserved $\mathbf{U}$ as follows:
+$(\mathbf{X},\mathbf{U})\mid A,G=E\sim\mathcal{N}([\frac{2A-1}{2},0],\begin{bmatrix}1&0\\ 0&1\end{bmatrix})$ (12)
+$(\mathbf{X},\mathbf{U})\mid A,G=O\sim\mathcal{N}([\frac{1-2A}{2},0],\begin{bmatrix}1&A-0.5\\ A-0.5&1\end{bmatrix}).$
+Finally, we generate the $T$-step short-term outcomes $\mathbf{S}$ and the long-term outcome $Y=S_{T+\mu}$ satisfying Assumption 4.1 as follows:
+$S_{t}(A)=A+0.1A\mathbf{X}+\mathbf{X}+\mathbf{U}+\sum_{k=0}^{t}S_{k}(A)+\epsilon_{S_{t}(A)}$ (13)
+where $\epsilon_{S_{t}(A)}$ are Gaussian noises. This will result in non-equal confounding bias in different time steps $t$, i.e., $\omega_{t}(\mathbf{x})=-2^{t}\mathbf{x}$.
+Baselines and implementation details We compare our method with the following baselines:
+  • LTEE [9].
+LTEE assumes the unconfoundedness assumption, and it uses a dual-headed RNN to capture the relationships of time-dependent surrogates and then estimates the effects using the predicted long-term outcomes.
+We use the code available at https://github.com/GitHubLuCheng/LTEE.
+  • Latent Unconfoundedness [5].
+Athey et al.
+[5] propose a long-term effect estimator under the latent unconfoundedness assumption.
+We implement this method following Sec.
+4.1 in [5] using correctly specified models.
+  • Sequential Outcomes [19].
+Imbens et al.
+[19] propose a long-term effect estimator by treating temporal short-term outcomes as proxies.
+we test three grouping strategies: (1) $S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$, (2) $S_{1},S_{\lceil T/2\rceil},S_{T}$, and (3) $S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$.
+We use the code available at https://github.com/CausalML/LongTermCausalInference/tree/main.
+  • NNPIV [28].
+Meza and Singh [28] propose an estimator for nested nonparametric instrumental variable regression, which is then applied to estimate the long-term heterogeneous effect under latent unconfoundedness and surrogacy assumption.
+We denote the estimators as NNPIV_LU and NNPIV_SInd under the latent unconfoundedness assumption [5] and surrogacy assumption [4], respectively.
+We use the code available at https://github.com/isaacmeza/NNPIV.
+  • CAECB [14].
+We implement CAECB [14] following the identification result in Eq.
+4.
+Since this estimator requires only using one short-term outcome, we implement four versions, including using the first, middle, last, and random one short-term outcome among all short-term outcomes.
+  • T-learner [26] trained using observational $Y$. This estimator is biased due to the latent confounding. We implement it using linear regression.
+  • T-learner [26] trained using experimental $Y$.
+We denote this estimator as $\hat{\tau}_{exp}$.
+This estimator is idealized and infeasible in practice since $Y$ is missing in the experimental data and serves as an idealized benchmark.
+We implement it using linear regression.
+  • Ours $\hat{\tau}$.
+We implement our method according to the algorithm described in Sec.
+5.1 using correctly specified models.
+Unless otherwise specified, we train the model using all short-term outcomes and omit step S0.
+All experiments are conducted on a machine equipped with an Intel(R) Xeon(R) E5-2620 v4 @ 2.10GHz CPU and a GeForce RTX 2080 Ti GPU. Each reported result is averaged over 50 runs.
+Full Results of Table 1.
+We present the full result of Table 1 in Tables 4 and 3.
+The results are expected and consistent with the analysis presented in the main text.
+Our estimator outperforms all baselines and closely aligns with the idealized estimator due to the correct assumption.
+Table 3: The results on synthetic data are presented as the mean and standard deviation.
+Lower values indicate better performance.
+The best-performing methods are bolded, with the exception of the idealized estimator $\hat{\tau}_{exp}$, which is excluded from the comparison.
+Data $\mu=1$, $T=6$ $\mu=2$, $T=6$ $\mu=3$, $T=6$ $\mu=4$, $T=6$ $\mu=5$, $T=6$
+Metrics ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$
+$\hat{\tau}_{exp}$ (idealized) 3.0033 $\pm$ 1.8226 2.0722 $\pm$ 1.7681 6.0076 $\pm$ 3.6446 4.1457 $\pm$ 3.5353 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 24.0297 $\pm$ 14.5790 16.5806 $\pm$ 14.1423 48.0590 $\pm$ 29.1580 33.1620 $\pm$ 28.2845
+T-learner (using obs.
+$Y$) 71.8431 $\pm$ 3.1612 6.8211 $\pm$ 2.8357 143.6844 $\pm$ 6.3217 13.6400 $\pm$ 5.6710 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3424 574.7384 $\pm$ 25.2877 54.5614 $\pm$ 22.6837 1149.4769 $\pm$ 50.5746 109.1205 $\pm$ 45.3685
+LTEE 73.0421 $\pm$ 7.1531 10.0977 $\pm$ 6.1071 142.1183 $\pm$ 10.9446 17.3357 $\pm$ 10.0487 287.3264 $\pm$ 28.5202 32.4417 $\pm$ 19.5739 589.0638 $\pm$ 56.6554 79.3176 $\pm$ 41.2797 1168.7612 $\pm$ 122.7687 137.0583 $\pm$ 107.5567
+Latent Unconfoundedness - 2.6322 $\pm$ 2.0120 - 5.2648 $\pm$ 4.0238 - 10.5293 $\pm$ 8.0471 - 21.0585 $\pm$ 16.0946 - 42.1179 $\pm$ 32.1874
+Sequential Outcomes ($S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$) - 67.5881 $\pm$ 3.0107 - 135.1792 $\pm$ 6.0220 - 270.3607 $\pm$ 12.0428 - 540.7164 $\pm$ 24.0884 - 1081.4361 $\pm$ 48.1759
+Sequential Outcomes ($S_{1},S_{\lceil T/2\rceil},S_{T}$) - 67.5790 $\pm$ 3.0138 - 135.1608 $\pm$ 6.0282 - 270.3237 $\pm$ 12.0553 - 540.6429 $\pm$ 24.1133 - 1081.2884 $\pm$ 48.2252
+Sequential Outcomes ($S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$) - 67.5797 $\pm$ 3.0151 - 135.1622 $\pm$ 6.0309 - 270.3266 $\pm$ 12.0607 - 540.6486 $\pm$ 24.1242 - 1081.3001 $\pm$ 48.2472
+NNPIV_LU 70.7124 $\pm$ 2.5320 6.1302 $\pm$ 2.8888 141.4309 $\pm$ 5.0655 12.2646 $\pm$ 5.7776 282.8595 $\pm$ 10.1320 24.5293 $\pm$ 11.5558 565.7158 $\pm$ 20.2586 49.0542 $\pm$ 23.1130 1131.4331 $\pm$ 40.5209 98.1118 $\pm$ 46.2263
+NNPIV_SInd 70.7155 $\pm$ 2.5329 6.1306 $\pm$ 2.8884 141.4320 $\pm$ 5.0665 12.2626 $\pm$ 5.7782 282.8655 $\pm$ 10.1314 24.5240 $\pm$ 11.5558 565.7307 $\pm$ 20.2627 49.0492 $\pm$ 23.1115 1131.4602 $\pm$ 40.5252 98.0989 $\pm$ 46.2232
+CAECB (using first $S$) 70.7257 $\pm$ 3.1153 6.7160 $\pm$ 2.7984 142.5669 $\pm$ 6.2757 13.5348 $\pm$ 5.6335 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 573.6209 $\pm$ 25.2417 54.4563 $\pm$ 22.6461 1148.3595 $\pm$ 50.5286 109.0154 $\pm$ 45.3308
+CAECB (using middle $S$) 67.3754 $\pm$ 2.9827 6.3982 $\pm$ 2.6884 139.2165 $\pm$ 6.1417 13.2171 $\pm$ 5.5210 282.9014 $\pm$ 12.4626 26.8566 $\pm$ 11.1910 570.2704 $\pm$ 25.1065 54.1385 $\pm$ 22.5317 1145.0089 $\pm$ 50.3933 108.6976 $\pm$ 45.2162
+CAECB (using last $S$) 36.1202 $\pm$ 2.0211 3.5778 $\pm$ 1.8603 107.9443 $\pm$ 4.9841 10.2576 $\pm$ 4.6541 251.6244 $\pm$ 11.2383 23.8972 $\pm$ 10.2153 538.9915 $\pm$ 23.8540 51.1791 $\pm$ 21.5097 1113.7291 $\pm$ 49.1287 105.7381 $\pm$ 44.1729
+CAECB (using random $S$) 59.8772 $\pm$ 11.9929 5.7194 $\pm$ 2.6871 131.7174 $\pm$ 12.7309 12.5383 $\pm$ 5.3841 275.4020 $\pm$ 16.0775 26.1778 $\pm$ 10.9897 562.7700 $\pm$ 26.2888 53.4597 $\pm$ 22.2995 1137.5094 $\pm$ 50.1542 108.0188 $\pm$ 44.9689
+Ours $\hat{\tau}$ 3.1110 $\pm$ 1.8230 2.1470 $\pm$ 1.7826 6.5594 $\pm$ 3.5589 4.4610 $\pm$ 3.5302 13.9945 $\pm$ 7.1682 9.2961 $\pm$ 7.0820 29.9866 $\pm$ 15.0884 19.3927 $\pm$ 14.4300 64.3076 $\pm$ 32.9868 40.5425 $\pm$ 29.7264
+Table 4: The results on synthetic data are presented as the mean and standard deviation.
+Lower values indicate better performance.
+The best-performing methods are bolded, with the exception of the idealized estimator $\hat{\tau}_{exp}$, which is excluded from the comparison.
+Data $\mu=5$, $T=4$ $\mu=4$, $T=5$ $\mu=3$, $T=6$ $\mu=2$, $T=7$ $\mu=1$, $T=8$
+Metrics ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$ ${\varepsilon_{PEHE}}$ $\varepsilon_{ATE}$
+$\hat{\tau}_{exp}$ (idealized) 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709 12.0142 $\pm$ 7.2895 8.2901 $\pm$ 7.0709
+T-learner (using obs.
+$Y$) 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423 287.3693 $\pm$ 12.6434 27.2795 $\pm$ 11.3423
+LTEE 296.1849 $\pm$ 27.7377 32.9559 $\pm$ 19.0871 288.1059 $\pm$ 31.1373 34.7556 $\pm$ 20.5223 287.3264 $\pm$ 28.5202 32.4417 $\pm$ 19.5739 289.8522 $\pm$ 25.2498 41.9068 $\pm$ 25.6233 284.0581 $\pm$ 26.0439 29.2638 $\pm$ 20.0548
+Latent Unconfoundedness - 10.5248 $\pm$ 8.0405 - 10.5302 $\pm$ 8.0423 - 10.5293 $\pm$ 8.0471 - 10.5270 $\pm$ 8.0476 - 10.5270 $\pm$ 8.0465
+Sequential Outcomes ($S_{\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1}$) - 270.3555 $\pm$ 12.0474 - 270.3607 $\pm$ 12.0428 - 270.3607 $\pm$ 12.0428 - 270.3272 $\pm$ 12.0610 - 270.3272 $\pm$ 12.0610
+Sequential Outcomes ($S_{1},S_{\lceil T/2\rceil},S_{T}$) - 270.3658 $\pm$ 12.0486 - 270.3213 $\pm$ 12.0654 - 270.3237 $\pm$ 12.0553 - 270.3323 $\pm$ 12.0565 - 270.3247 $\pm$ 12.0560
+Sequential Outcomes ($S_{1:\lceil T/2\rceil-1},S_{\lceil T/2\rceil},S_{\lceil T/2\rceil+1:T}$) - 270.3554 $\pm$ 12.0473 - 270.3266 $\pm$ 12.0607 - 270.3266 $\pm$ 12.0607 - 270.3295 $\pm$ 12.0557 - 270.3295 $\pm$ 12.0557
+NNPIV_LU 282.8435 $\pm$ 10.1334 24.6196 $\pm$ 11.5648 282.8424 $\pm$ 10.1261 24.5424 $\pm$ 11.5495 282.8595 $\pm$ 10.1320 24.5293 $\pm$ 11.5558 282.8690 $\pm$ 10.1315 24.5386 $\pm$ 11.5653 282.8671 $\pm$ 10.1264 24.5355 $\pm$ 11.5659
+NNPIV_SInd 282.8679 $\pm$ 10.1239 24.5522 $\pm$ 11.5413 282.8408 $\pm$ 10.1337 24.5238 $\pm$ 11.5430 282.8655 $\pm$ 10.1314 24.5240 $\pm$ 11.5558 282.8683 $\pm$ 10.1288 24.5331 $\pm$ 11.5663 282.8679 $\pm$ 10.1240 24.5326 $\pm$ 11.5657
+CAECB (using first $S$) 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 286.2518 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 286.2519 $\pm$ 12.5973 27.1744 $\pm$ 11.3048 286.2519 $\pm$ 12.5973 27.1744 $\pm$ 11.3048
+CAECB (using middle $S$) 285.1353 $\pm$ 12.5525 27.0677 $\pm$ 11.2662 282.9014 $\pm$ 12.4625 26.8566 $\pm$ 11.1910 282.9014 $\pm$ 12.4626 26.8566 $\pm$ 11.1910 278.4314 $\pm$ 12.2812 26.4329 $\pm$ 11.0441 278.4314 $\pm$ 12.2812 26.4330 $\pm$ 11.0441
+CAECB (using last $S$) 278.4314 $\pm$ 12.2812 26.4330 $\pm$ 11.0441 269.4957 $\pm$ 11.9268 25.5886 $\pm$ 10.7549 251.6244 $\pm$ 11.2383 23.8972 $\pm$ 10.2153 215.8874 $\pm$ 9.9675 20.5128 $\pm$ 9.3073 144.4755 $\pm$ 8.0850 14.3059 $\pm$ 7.4382
+CAECB (using random $S$) 283.3244 $\pm$ 13.2920 26.8862 $\pm$ 11.1734 280.1459 $\pm$ 13.3821 26.5476 $\pm$ 11.0112 275.4020 $\pm$ 16.0775 26.1778 $\pm$ 10.9897 270.3812 $\pm$ 22.4630 25.5679 $\pm$ 10.7551 254.5652 $\pm$ 45.7734 24.0325 $\pm$ 10.6828
+Ours $\hat{\tau}$ 19.1631 $\pm$ 12.1030 10.3532 $\pm$ 9.2908 16.6761 $\pm$ 13.1391 10.4123 $\pm$ 9.5082 13.9945 $\pm$ 7.1682 9.2961 $\pm$ 7.0820 12.7653 $\pm$ 7.4767 8.5308 $\pm$ 7.0285 12.3040 $\pm$ 7.3958 8.5348 $\pm$ 7.1581
+Appendix E Proof of Theorem 3.7
 We first restate Theorem 3.7 as follows:
-Theorem A.1.
+Theorem E.1.
 Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 3.6 hold, then $\tau(\mathbf{x})$ can be identified as follows:
 $\tau(\mathbf{x})$ (14)
-$=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$
+$=$ $\mathbb{E}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]$
 $=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+\mu_{S}^{E}(1,\mathbf{x})-\mu_{S}^{E}(0,\mathbf{x})+\mu_{S}^{O}(0,\mathbf{x})-\mu_{S}^{O}(1,\mathbf{x}).$
@@
 $\tau(\mathbf{X})$ (15)
-$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$
-$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
-$+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
-$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]$
-+$+\{\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]\}\times p(A=0|\mathbf{X},G=O)$
-+$+\{\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]\}\times p(A=1|\mathbf{X},G=O)$
-$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
-+$+\{\mathbb{E}\left[S(1)|\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(1)|\mathbf{X},G=O,A=1\right]\}\times p(A=0|\mathbf{X},G=O)$
-+$+\{\mathbb{E}\left[S(0)|\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(0)|\mathbf{X},G=O,A=1\right]\}\times p(A=1|\mathbf{X},G=O),$
+$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O]$
+$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)-\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)$
+$+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)$
+$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]$
++$+\{\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]\}\times p(A=0\mid\mathbf{X},G=O)$
++$+\{\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]\}\times p(A=1\mid\mathbf{X},G=O)$
+$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
++$+\{\mathbb{E}\left[S(1)\mid\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(1)\mid\mathbf{X},G=O,A=1\right]\}\times p(A=0\mid\mathbf{X},G=O)$
++$+\{\mathbb{E}\left[S(0)\mid\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(0)\mid\mathbf{X},G=O,A=1\right]\}\times p(A=1\mid\mathbf{X},G=O),$
 where the first equality is based on Assumption 3.5 and the last equality is based on Assumption 3.6. Similarly, for short-term conditional causal effects, we have:
-$\mathbb{E}[S(1)|\mathbf{X},G=O]-\mathbb{E}[S(0)|\mathbf{X},G=O]$ (16)
-$=$ $\mathbb{E}[S|\mathbf{X},G=O,A=1]-\mathbb{E}[S|\mathbf{X},G=O,A=0]$
-+$+\{\mathbb{E}[S(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S(1)|\mathbf{X},G=O,A=1]\}\times p(A=0|\mathbf{X},G=O)$
-+$+\{\mathbb{E}[S(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S(0)|\mathbf{X},G=O,A=1]\}\times p(A=1|\mathbf{X},G=O)$
+$\mathbb{E}[S(1)\mid\mathbf{X},G=O]-\mathbb{E}[S(0)\mid\mathbf{X},G=O]$ (16)
+$=$ $\mathbb{E}[S\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S\mid\mathbf{X},G=O,A=0]$
++$+\{\mathbb{E}[S(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S(1)\mid\mathbf{X},G=O,A=1]\}\times p(A=0\mid\mathbf{X},G=O)$
++$+\{\mathbb{E}[S(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S(0)\mid\mathbf{X},G=O,A=1]\}\times p(A=1\mid\mathbf{X},G=O)$
 Then, combining Eq. (15) and (16), we have
 $\tau(\mathbf{X})$ (17)
-$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$
-$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
-$+\mathbb{E}[S(1)|\mathbf{X},G=O]-\mathbb{E}[S(0)|\mathbf{X},G=O]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
-$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
-$+\mathbb{E}[S(1)|\mathbf{X},G=E]-\mathbb{E}[S(0)|\mathbf{X},G=E]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
-$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
-$+\mathbb{E}[S|\mathbf{X},G=E,A=1]-\mathbb{E}[S|\mathbf{X},G=E,A=0]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
+$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O]$
+$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
+$+\mathbb{E}[S(1)\mid\mathbf{X},G=O]-\mathbb{E}[S(0)\mid\mathbf{X},G=O]-\mathbb{E}[S\mid\mathbf{X},G=O,A=1]+\mathbb{E}[S\mid\mathbf{X},G=O,A=0]$
+$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
+$+\mathbb{E}[S(1)\mid\mathbf{X},G=E]-\mathbb{E}[S(0)\mid\mathbf{X},G=E]-\mathbb{E}[S\mid\mathbf{X},G=O,A=1]+\mathbb{E}[S\mid\mathbf{X},G=O,A=0]$
+$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
+$+\mathbb{E}[S\mid\mathbf{X},G=E,A=1]-\mathbb{E}[S\mid\mathbf{X},G=E,A=0]-\mathbb{E}[S\mid\mathbf{X},G=O,A=1]+\mathbb{E}[S\mid\mathbf{X},G=O,A=0]$
 $=$ $\mu_{Y}^{O}(1,\mathbf{X})-\mu_{Y}^{O}(0,\mathbf{X})+\mu_{S}^{E}(1,\mathbf{X})-\mu_{S}^{E}(0,\mathbf{X})+\mu_{S}^{O}(0,\mathbf{X})-\mu_{S}^{O}(1,\mathbf{X}),$
 where the second equality is based on Assumption 3.5 and the last equality is based on Assumption 3.4. This finishes our proofs. ∎
-Appendix B Proof of Proposition 4.3
+Appendix F Proof of Proposition 4.3
 We first restate the proposition as follow:
-Proposition B.1.
+Proposition F.1.
 Under Assumption 4.1, $\forall t$ the confounding biases between times $t$ and $t+1$ follow
@@
 We start from the long-term causal effects at time step $t$ in experimental data $G=E$:
-$\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]$ (19)
-=$\overset{(a)}{=}$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O]$
-$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
-$+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
-$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]$
-+$+\{\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
-+$+\{\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
-$=$ $\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
+$\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=E]$ (19)
+=$\overset{(a)}{=}$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O]$
+$=$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$=$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$+\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)-\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)$
+$+\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)$
+$=$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]$
++$+\{\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
++$+\{\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
+$=$ $\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=0]$
 $+b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
 where the equality $(a)$ is based on Assumption 3.5. By rewriting the last equality above, we obtain:
-$\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]$ (20)
-$=$ $\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
+$\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=E]$ (20)
+$=$ $\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=0]$
 $+b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
-$\iff$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]+\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
+$\iff$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=E]-\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=1]+\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=0]$
 $=$ $b_{t}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{t}(0,\mathbf{X})p(A=1,\mathbf{X},G=O),$
@@
 where $(a)$ is based on Assumption 4.1, i.e., $b_{t+1}(a,\mathbf{X})=f(\mathbf{X})b_{t}(a,\mathbf{X})$, and $(b)$ is based on Eq. (20) and finishes our proof. ∎
-Appendix C Proof of Theorem 4.5
+Appendix G Proof of Theorem 4.7
 We first restate Theorem 4.7 as follows:
-Theorem C.1.
+Theorem G.1.
 Suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5 and 4.1 hold, then $\tau(\mathbf{x})$ can be identified as follows:
 $\tau(\mathbf{x})$ (22)
-$=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$
+$=$ $\mathbb{E}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]$
 $=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{X})\left(\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})\right).$
 Proof.
-$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$ (23)
-$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
-$+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
-$+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
-$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]$
-+$+\{\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
-+$+\{\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
-$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
+$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O]$ (23)
+$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
+$+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)-\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)$
+$+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)$
+$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]$
++$+\{\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
++$+\{\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
+$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
 $+b_{T+\mu}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{T+\mu}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)$
-where the first equality is based on Assumption 3.5, i.e., $G{\perp\!\!\!\perp}Y(a)|\mathbf{X}$, and last equality is based on Assumption 4.1.
+where the first equality is based on Assumption 3.5, i.e., $G{\perp\!\!\!\perp}Y(a)\mid\mathbf{X}$, and last equality is based on Assumption 4.1.
 Further based on Assumption 4.1, we have $b_{T+\mu}(a,\mathbf{X})=f^{\mu}(\mathbf{X})b_{T}(a,\mathbf{X})$. Then, we rewrite the equality above as:
-$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$ (24)
-$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
+$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O]$ (24)
+$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
 $+f^{\mu}(\mathbf{X})b_{T}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-f^{\mu}(\mathbf{X})b_{T}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)$
 Similarly, for short-term ITE at last time step $T$ we have:
-$\mathbb{E}[S_{T}(1)|\mathbf{X}]-\mathbb{E}[S_{T}(0)|\mathbf{X}]$ (25)
-$=$ $\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
-+$+\{\mathbb{E}[S_{T}(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
-+$+\{\mathbb{E}[S_{T}(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
-$=$ $\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
+$\mathbb{E}[S_{T}(1)\mid\mathbf{X}]-\mathbb{E}[S_{T}(0)\mid\mathbf{X}]$ (25)
+$=$ $\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=0]$
++$+\{\mathbb{E}[S_{T}(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(1)\mid\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
++$+\{\mathbb{E}[S_{T}(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(0)\mid\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
+$=$ $\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=0]$
 $+b_{T}(1,\mathbf{X})p(A=0,\mathbf{X},G=O)-b_{T}(0,\mathbf{X})p(A=1,\mathbf{X},G=O)$
 Then, combining Eq. 24 and 25, we have
-$\tau(\mathbf{X})=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(1)|\mathbf{X}]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(0)|\mathbf{X}]$ (26)
-$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
-$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=E,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=E,A=0]$
-$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
+$\tau(\mathbf{X})$ (26)
+$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(1)\mid\mathbf{X}]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(0)\mid\mathbf{X}]$
+$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=0]$
+$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=E,A=1]$
+$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=E,A=0]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=0]$
 $=$ $\mu_{Y}^{O}(1,\mathbf{x})-\mu_{Y}^{O}(0,\mathbf{x})+f^{\mu}(\mathbf{X})\left(\mu_{S_{T}}^{E}(1,\mathbf{x})-\mu_{S_{T}}^{E}(0,\mathbf{x})+\mu_{S_{T}}^{O}(0,\mathbf{x})-\mu_{S_{T}}^{O}(1,\mathbf{x})\right)$
 where the second equality is based on Assumption 3.5. This finishes our proof. ∎
-Appendix D Proof of Lemma 5.4
+Appendix H Proof of Lemma 5.4
 We first state the following lemma that is used in the proof of Lemma 5.4.
-Lemma D.1.
+Lemma H.1.
 suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1 hold, then $\forall t$,
@@
 ∎
-We now restate our Lemma 5.4 as follows and formally prove that.
-Lemma D.2.
-Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, 5.2 and 5.3 hold, then we have
-$\hat{f}(\mathbf{x})-f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$ (29)
+We now restate our Lemma 5.4 as follows.
+Lemma H.2.
+Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, 5.2, and 5.3 hold, then we have
+$\hat{f}(\mathbf{x})-$ $f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right),$ (29)
+which attains the oracle rate if $\min\{\alpha,\beta\}\geq\eta$.
 Proof.
-We apply Proposition 1 of Kennedy et al. (Kennedy, 2023), yielding that
+We apply Proposition 1 in Kennedy et al. [25], yielding that
 $\hat{f}(\mathbf{x})-f(\mathbf{x})=$ $(\hat{f}(\mathbf{x})-\tilde{f}(\mathbf{x}))+(\tilde{f}(\mathbf{x})-f(\mathbf{x}))$ (30)
 $=$ $(\hat{f}(\mathbf{x})-\tilde{f}(\mathbf{x}))+O_{p}(R_{n}^{*}(\mathbf{x}))$
-$=$ $\hat{\mathbb{E}}_{n}[\hat{r}(\mathbf{X})|\mathbf{X}=\mathbf{x}]+o_{p}(R_{n}^{*}(\mathbf{x}))+O_{p}(R_{n}^{*}(\mathbf{x}))$
-where $\hat{r}(\mathbf{x})=\mathbb{E}[\hat{f}(\mathbf{X})|\mathbf{X}=\mathbf{x}]-f(\mathbf{x})$, and $\mathcal{R}_{n}^{*}(x)$ is the oracle risk of second-stage regression and further under Assumption 5.2, we know $f$ is $\eta$-smooth, thus $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$.
+$=$ $\hat{\mathbb{E}}_{n}[\hat{r}(\mathbf{X})\mid\mathbf{X}=\mathbf{x}]+o_{p}(R_{n}^{*}(\mathbf{x}))+O_{p}(R_{n}^{*}(\mathbf{x}))$
+where $\hat{r}(\mathbf{x})=\mathbb{E}[\hat{f}(\mathbf{X})\mid\mathbf{X}=\mathbf{x}]-f(\mathbf{x})$, and $\mathcal{R}_{n}^{*}(x)$ is the oracle risk of second-stage regression and further under Assumption 5.2, we know $f$ is $\eta$-smooth, thus $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$.
 Here, $f$ is a time-series model ($1$-order autoregressive), optimized with $T-1$-length steps and $n$ samples, as shown in Eq.
@@
 where the inequality $(a)$ is based on $(a+b)^{2}\leq 2(a^{2}+b^{2})$, and $(b)$ is based on Assumption 5.3.
-Under Assumption 5.3, $\mu_{S_{t}}^{O}$, $\mu_{S_{t}}^{E}$ and their estimates are all bounded, and thus $\hat{\omega}_{t}(\mathbf{x})$ and its estimates are also bounded, thus $(b)$ holds.
-Then, applying Lemma D.1, we know for a fixed time step $t$, $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}-\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}=O_{p}(n^{\frac{-\alpha}{2\alpha+d}}+n^{\frac{-\beta}{2\beta+d}})$.
+Under Assumption 5.3, $\mu_{S_{t}}^{O}$, $\mu_{S_{t}}^{E}$ and their estimates are all bounded, and thus $\omega_{t}(\mathbf{x})$ and its estimates are also bounded, thus $(b)$ holds.
+Then, applying Lemma H.1, we know for a fixed time step $t$, $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}-\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}=O_{p}(n^{\frac{-\alpha}{2\alpha+d}}+n^{\frac{-\beta}{2\beta+d}})$.
 Then, for $\hat{r}(\mathbf{x})$, we have the effective sample is of size $(T-1)n$, and then combining with Eq.
 (30) and $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$, we obtain the desired result.
+And attaining the oracle rate requires $(\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}\geq\max\{(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}},(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\}$, yielding $\min\{\alpha,\beta\}\geq\eta$.
 ∎
-Appendix E Proof of Theorem 5.5
+Appendix I Proof of Theorem 5.5
 We first restate Theorem 5.5 as follows:
-Theorem E.1.
+Theorem I.1.
 Suppose Lemma 5.4 hold, then we have
-$\hat{\tau}(\mathbf{x})-\tau(\mathbf{x})=O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$ (32)
+$\hat{\tau}(\mathbf{x})-\tau(\mathbf{x})$ (32)
+$=$ $O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$
 Proof.